### v2

# Training PINNs with Hard Constraints and Adaptive Weights: An Ablation Study

Abstract
Training Physics-Informed Neural Networks (PINNs) to solve stiff time-dependent partial differential equations (PDEs) remains a significant challenge. The main difficulties lie in precisely enforcing initial conditions and balancing the various loss components. In stiff PDEs, where solutions show rapid transitions or sharp gradients, this balance becomes especially difficult. The wide range of scales in the dynamics can cause certain losses to dominate, leading to unstable or inefficient training. This paper presents a comprehensive ablation study focused on two pivotal training schemes: the enforcement of hard constraints for initial and boundary conditions, and the implementation of adaptive weights. We specifically examine their impact on the performance of PINNs when applied to stiff time-dependent PDEs from materials science and mathematical biology applications. We conduct extensive numerical experiments across a diverse range of time-dependent PDEs from Allen-Cahn, Cahn-Hillard, to Gray-Scott systems. We further discuss the implications of our findings for improving the robustness and efficiency of PINN training, particularly in settings where accurate representation of initial conditions and balanced loss contributions are paramount.
1 Introduction
Time-dependent partial differential equations (PDEs) constitute the mathematical foundation of numerous scientific and engineering disciplines, serving as essential tools for modeling a diverse range of physical phenomena. Their applications span from the intricate dynamics of fluid flow, as described by the Navier–Stokes equations, to the complex interplay of chemical reactions in biological systems, captured by reaction–diffusion models. Time-dependent PDEs offer a powerful framework for understanding and predicting the evolution of systems over time whitham2011linear ; burgers1948mathematical ; allen1979microscopic ; shen2010numerical ; bazant2017thermodynamic ; horstmann2013precipitation ; tian2015electrochemically ; hyman2014liquid ; takatori2015towards ; cahn1958free ; lee2014physical ; miranville2017cahn ; kudryashov1990exact ; michelson1986steady ; de1991turing ; nishiura1999skeleton ; zhabotinsky2007belousov ; cassani2021belousov ; kato1987nonlinear ; kevrekidis2001discrete ; fibich2015nonlinear . In particular, the imposition of periodic boundary conditions allows for the simulation of phenomena over effectively infinite physical domains, providing valuable insights into the intrinsic behavior of physical systems. However, even the numerical solution of these PDEs, especially in scenarios involving stiffness or complex non-linearities, can be exceedingly challenging, often demanding significant computational resources and specialized numerical techniques.
The advent of machine learning has ushered in a new era of scientific computing, offering powerful tools for data-driven modeling and analysis. Machine learning methodologies have shown remarkable success in various science/engineering fields, revolutionizing areas such as protein folding prediction PF2019 , drug discovery and development CARRACEDOREBOREDO20214538 , and even optimizing fundamental algorithmic processes such as matrix-vector multiplication MMML2022 . However, traditional machine learning approaches, which rely on vast datasets, can prove impractical in scientific contexts where data acquisition is expensive, time consuming, or simply infeasible. Furthermore, purely data-driven models often lack physical fidelity and model interpretability, as they may fail to capture the underlying physical principles governing the systems they aim to describe. Physics-informed machine learning (PIML), specifically physics-informed neural networks (PINNs), address these limitations by seamlessly integrating physical laws into the learning process, significantly reducing the amount of training data needed. By embedding constraints derived from the governing PDEs directly into the neural network architecture, PINNs enable accurate predictions even with sparse and noisy data yang2019physics ; chen2021solving ; raissi2018hidden ; raissi2019physics ; liu2019multi ; liu2021dual ; jin2021nsfnets ; cai2021deepm ; rad2020theory ; zhu2021machine ; chen2022using ; wang2023pinnsbased . Despite the promising applications, effective training of PINNs, especially for stiff time-dependent PDEs, remains a significant challenge. The inherent difficulty in propagating information from initial and boundary data points into the computational domain, coupled with the need to balance multiple loss terms, often leads to convergence issues and suboptimal solutions wang2020understanding ; wang2020and ; mcclenny2020self ; coutinho2022physics .
This paper presents an ablation study of training algorithms for PINNs applied to stiff partial differential equations (PDEs), focusing specifically on two schemes: hard constraints and self-adaptive loss weights mcclenny2020self . The hard-constraint scheme directly embeds information about initial and boundary conditions into the neural network architecture. This transformation reduces the effects of spectral bias from training various kinds of loss terms, leading to improved solution accuracy wang2020and . PINN’s training can treated as a regularized regression, where initial and boundary conditions, along with observational data, serve as training data, and the physics-informed loss acts as regularization. However, the multi-objective nature of the loss function and the sensitivity of PDE solutions to initial conditions can significantly hinder the search for global minimizers. By explicitly incorporating boundary and initial condition data into the network transformation, we streamline the training process, particularly for stiff time-dependent PDEs, where accurately capturing the initial conditions is critical. In addition, we investigate the use of adaptive loss weights to alleviate the spectral bias from the multi-objective loss in PINN training mcclenny2020self . Through our ablation study of these training schemes, we arrive at the conclusion that hard constrain transformation provides better improvement. We emphasize that hard constraints and self-adaptive weights are not the only effective strategies for handling stiff PDEs. These methods can be integrated with other advanced PINN techniques, such as causal PINNs wang2020and , time-marching PINNs wight2020solving , residual-based-attention (RBA) PINNs ANAGNOSTOPOULOS2024116805 , curriculum training krishnapriyan2021characterizing , and co-training PINNs zhong2024label , to achieve synergistic improvements in prediction accuracy and robustness. Through a rigorous ablation study, we demonstrate the efficacy of the proposed approach across a range of stiff time-dependent PDEs, offering valuable insights into optimal training strategies for PINNs in these challenging scenarios.
1.1 Related Work
PINNs have shown significant potential to solve a wide range of PDEs. However, challenges arise, particularly when dealing with “stiff” PDEs, where explicit numerical methods require extremely small time steps for stability. A primary concern is the imbalance between the fitting for initial data and residual components within the PINN loss function wight2020solving ; shin2020convergence ; wang2020understanding ; wang2020and . This imbalance often leads to a gradient descent that favors minimization of residual loss over data fitting loss, hindering convergence to accurate solutions. This issue is exacerbated in forward problems, where data is primarily concentrated at initial and boundary conditions, leaving limited information within the PDE domain. Specifically, in time-dependent problems, PINNs struggle to propagate information from initial conditions to subsequent time steps, a phenomenon widely documented in the literature wight2020solving ; mcclenny2020self ; krishnapriyan2021characterizing ; wang2022respecting ; haitsiukevich2022improved .
Several strategies have been proposed to address these challenges. One approach, introduced by Wight and Zhao wight2020solving , involves time-marching, where the time domain is divided into smaller intervals, and PINNs are sequentially trained on each interval, starting from the initial condition. While effective, this method can be computationally expensive due to the need to train multiple networks. To improve information propagation, researchers have emphasized the importance of weighting initial condition data and residual points near the initial time wight2020solving ; mcclenny2020self . McClenny and Konstantinidis mcclenny2020self demonstrated that their self-adaptive PINNs can autonomously learn these weights during training, particularly in scenarios with complex initial conditions, such as the wave equation.
Further advancements include the “sequence-to-sequence” approach proposed by Krishnapriyan et al. krishnapriyan2021characterizing , which sequentially learns each time step in a time-marching scheme. Wang et al. wang2022respecting identified the lack of “causality” in standard PINN training and proposed a corrective weighting scheme that gradually prioritizes data from early to later times, leveraging information from the PDE residue. This approach shares similarities with the self-adaptive weighting mechanism of McClenny and Konstantinidis mcclenny2020self . Additionally, Haitsiukevich et al. haitsiukevich2022improved introduced a pseudo-label approach, resembling self-training, to further enhance PINN convergence and accuracy. In zhong2024label , a mixed training scheme using labal-propagation between data points and collocation points is introduced and tested. For stiff parabolic PDEs such as viscous Burgers with small viscosity, the authors in coutinho2022physics proposed an data-driven artificial viscosity scheme to produce stable solutions and accurate to capture sharp transitions in the solutions.
Beyond loss weighting and time marching, structural modifications to PINNs have also been explored. Braga-Neto et al. braganeto2023characteristicsinformed introduced Characteristic-Informed Neural Networks (CINNs) for transport equations, where PDE information is integrated into the neural network architecture. A similar approach, albeit for Dirichlet boundary conditions, was proposed by Lagaris et al. Lagaris_1998 . Our method builds upon these architectural modifications, with a focus on enhancing stability and ensuring proper propagation of information, especially for stiff time-dependent PDEs. We carefully design our architecture to address the challenges of information propagation and loss balancing, aiming to improve the robustness and efficiency of PINN training.
2 Methodology
We consider the following setup for a family of time dependent PDEs. Let $u$ be an unknown function defined on $[0,T]\times\Omega$ with $\Omega\subset\mathcal{R}^{d}$. Here the physical domain $\Omega$ comes with a Lipschitz boundary $\partial\Omega$ and $\bar{\Omega}=\Omega\cup\partial\Omega$. Let $\mathcal{P}$ be an operator which can include differentiation and integration. For example, for the $1D$ heat equation $u_{t}-\lambda u_{xx}=0$, $\mathcal{P}=\lambda\frac{\partial^{2}}{\partial x^{2}}$. Notice that $\mathcal{P}$ can be nonlinear. Let $\mathcal{B}$ be another operator on the boundary, that is, for Dirichlet boundary condition $u(t,x)=g(t,x)$ for $(t,x)\in[0,T]\times\partial\Omega$, $\mathcal{B}=\mathcal{I}$ (the identity). Then we say that $u$ is a solution of an PDE, if $u$ satisfies the following
$\begin{cases}u_{t}-\mathcal{P}(u)(t,\mathbf{x})&=f(t,\mathbf{x}),\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ u(0,\mathbf{x})&=u_{0}(\mathbf{x}),\quad\mathbf{x}\in\bar{\Omega},\\ \mathcal{B}(u)(t,\mathbf{x})&=g(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega,\\ \end{cases}$ (1)
Here the functions $u_{0}$, $f$, and $g$ are user-inputs and are assumed to satisfy desired regularity so that the existence and uniqueness of solutions for such a PDE is guaranteed. We also assume that the compactibility condition, that is, $g(0,\mathbf{x})=u_{0}(\mathbf{x})$ for $\mathbf{x}\in\partial\Omega$, is satisfied to prevent any ill-conditioning of the PDE problem.
Types of Boundary Conditions: one can consider many kinds of boundary conditions as follows
$\mathcal{B}(u)(t,\mathbf{x})$ $=u(t,\mathbf{x}),\quad\text{Dirichlet type},$
$\mathcal{B}(u)(t,\mathbf{x})$ $=\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Neumann type, $\mathbf{n}$ is the outward normal vector to $\partial\Omega$},$
$\mathcal{B}(u)(t,\mathbf{x};a,b)$ $=au(t,\mathbf{x})+b\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Robin type}.$
The periodic boundary condition gives $d$ different equations, and they are given as follows,
$\mathcal{B}(u)(t,\mathbf{x})=u(t,\mathbf{x})-u(t,\mathbf{x}+P\mathbf{e}_{i})=0,\quad i=1,\cdots,d.$ (2)
Here $P$ is the period and $\mathbf{e}_{i}$ is the $i^{th}$ standard basis vector in $\mathcal{R}^{d}$, that is,
$\mathbf{e}_{i}=\underbrace{\begin{bmatrix}0&\cdots&0&1&0&\cdots&0\end{bmatrix}^{\top}}_{\text{$1$ is the in the $i^{th}$ position.}}.$
One can even consider a mixture of two or all of the above BC conditions. For example, we can have $\Omega=\Gamma_{1}\cup\cdots\cup\Gamma_{K}$ with each $\Gamma_{i}$ defines a different kind of BC. In this paper, we will mainly focus on the periodic type of boundary conditions, however other kinds of BCs can be
Classical numerical techniques such as the Finite Difference Method (which approximates derivatives on a point mesh), Finite Element Method (utilizing basis approximation on a triangular mesh with weak formulation), and Spectral Method (employing Fourier basis with FFT on a point mesh) are widely utilized in scientific computations. However, recent advancements in scientific machine learning integrate the principles of physics, particularly the understanding of PDEs, into the training process of machine learning models. This integration enables the development of PINNs to solve for $u$ as follows: one tries to find an approximate solution from a set of deep neural networks (of the same depth, same number of neurons on each hidden layer, and same activation functions on each layer) $\mathcal{H}_{\text{NN}}$, which is a minimizer of the following loss functional
$\text{Loss}(u_{nn})=\text{Data Loss}+\lambda*\text{Physical Loss}$
where $\lambda$ is a regularization parameter, and the two losses are as follows
Data $\text{Data Loss}=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
$\quad+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2}$
Physical $\text{Physical Loss}=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}$
for $u_{nn}\in\mathcal{H}_{\text{NN}}$. Here $\{(t^{CL},\mathbf{x}^{CL})_{i}\}_{i=1}^{N_{CL}}\in(0,T]\times\Omega$ are called collocation points, $\{(0,\mathbf{x}^{IC})_{i}\}_{i=1}^{N_{IC}}\in\{0\}\times\bar{\Omega}$ are the initial condition points, and $\{(t^{BC},\mathbf{x}^{BC})_{i}\}_{i=1}^{N_{BC}}\in[0,T]\times\partial\Omega$ are the boundary condition points. The minimizer, denoted as $u_{NN}$, will an be approximate solution to (1).
2.1 Hard Constrains
The main motivation for us to consider hard constrains transformation for PINNs is the possible ill-conditioning of the PDE solution map’s dependence on IC and BC. Traditional numerical methods do not need to encounter this kind of difficulty, as the IC data is exactly satisfied in their solutions. PINNs, on the other hand, use $L^{2}$-loss to fit the IC/BC and PDE residual data, which causes the solutions to be highly sensitive to the tightness of fit for the IC in the training (as pointed out in mcclenny2020self ). By transforming the need of fitting IC/BC conditions out of PINNs, it reduces the spectral bias in the training. Therefore, We consider the following transformation
$\tilde{u}(t,\mathbf{x})=\psi(t,\mathbf{x})+\phi(t,\mathbf{x})u_{nn}(t,\mathbf{x})$
Here the functions $\psi$ and $\phi$ are smooth functions with the following property
$\psi(0,\mathbf{x})=u_{0}(\mathbf{x}),\quad\phi(0,\mathbf{x})=0,\quad\text{for $\mathbf{x}\in\bar{\Omega}$},$
where in the case of periodic BC, we require $\psi$ and $\phi$ to be both $P$-periodic; for other BC types, we instead require $\phi$ and $\psi$ to satisfy
$\psi(t,\mathbf{x})=g(t,\mathbf{x}),\quad\phi(t,\mathbf{x})=0,\quad\text{for $(t,\mathbf{x})\in[0,T]\times\partial\Omega$}.$
When such transformation is used, then the training for $u_{nn}$ is changed, due to the fact that IC/BC can be built into the PDE loss, i.e.
$\text{Loss}(u_{nn})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\tilde{u}_{t}-\mathcal{P}(\tilde{u}))(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2},\quad\text{for $\tilde{u}=\psi+\phi u_{nn}$}.$
We have reduced the types of losses in the training; however the information about IC and BC has then moved into the PDE residual loss through $\psi$ and $\phi$. The difficulty in training the new PDE residual loss has been increased. Hence, other training enhancement techniques can be used to improve the new PDE residual training.
Periodic Neural Networks: even though we require $\psi$ and $\phi$ to be periodic, we still need to build the information of periodicity into $u_{nn}$, hence we consider the following composition
$u_{nn}(t,x)=f_{nn}(v(t,x)),\quad\text{$f_{nn}$ is a neural network, and $x\in\mathcal{R}$},$
and
$v(t,x)=\begin{bmatrix}t&1&\cos(\frac{2\pi}{P}x)&\sin(\frac{2\pi}{P}x)&\cos(\frac{2\pi}{P}2x)&\cdots&\cos(\frac{2\pi}{P}mx)&\sin(\frac{2\pi}{P}mx)\end{bmatrix}^{\top},$
where $m>0$ is a positive integer (another hyper-parameter). With this composition, $u_{nn}$ will automatically become $P$-periodic dong2021method . For higher dimensional physical $\mathbf{x}$, similar transformation can be used.
2.2 Mini-batching
We have eliminated the IC/BC losses, but the transformed PDE loss might create additional difficulties in training due to the introduction of IC/BC functions into the PDE. We adopt a machine learning technique, called mini-batching, commonly used to reduce training complexity by evaluating the loss on randomly selected subsets of collocation data. In PINNs, this reduces memory costs, especially when handling large numbers of collocation points. Sampling mini-batches from collocation points also helps stabilize and accelerate optimization. So we consider mini-batching in the training for decreasing the PDE residual loss. As pointed out in wight2020solving , the mini-batching technique is similar to time-marching resampling. But with mini-batching, one does not employ an explicit time grid.
2.3 Self-Adaptive Physics-Informed Neural Networks
The proper choice of penalty term $\lambda$ can be updated from derived formula as mentioned in wang2020and . However, data-driven discovery of $\lambda$ is possible. Outline in mcclenny2020self , the loss is further modified as
$\text{Data Loss}(\lambda_{IC,i}\}_{i=1}^{N_{IC}},\{\lambda_{BC,i}\}_{i=1}^{N_{BC}})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}m(\lambda_{IC,i})|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
$\quad+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}m(\lambda_{BC,i})|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2}$
Physical $\text{Physical Loss}(\{\lambda_{PDE,i}\}_{i=1}^{N_{PDE}})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}m(\lambda_{PDE,i})|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}$
The set of loss weights, $\{\lambda_{IC,i}\}_{i=1}^{N_{IC}}\cup\{\lambda_{BC,i}\}_{i=1}^{N_{BC}}\cup\{\lambda_{PDE,i}\}_{i=1}^{N_{PDE}}$, are designed in a way to increase when needed during any training epoch in order to combat the spectral bias. Adaptive methods are essential to ensure that the neural network accurately addresses the challenging spots in the solutions of “stiff” PDEs. In our experiments, we employ self-adaptive PINNs as proposed by mcclenny2020self to train PINNs adaptively. This approach involves fully trainable adaptation weights applied individually to each training point, allowing the neural network to independently identify and focus on the challenging regions of the solution.
2.4 Performance Measures
Let $\{x_{i},t_{i}\}_{i=1}^{N}$ be a collection of $N$ data points where we evaluate both the reference solution $u(x,t)$ and the neural network output $\mathcal{U}(x,t)$. To measure the accuracy of the trained model, we compute the relative $L_{1}$ norm $\mathcal{E}_{1}$ and relative $L_{2}$ norm $\mathcal{E}_{2}$, defined by:
$\mathcal{E}_{1}=\frac{\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|}{\sum_{i=1}^{N}|u(x_{i},t_{i})|},\quad\quad\quad\mathcal{E}_{2}=\frac{\Big{(}\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|^{2}\Big{)}^{1/2}}{\sqrt{\sum_{i=1}^{N}|u(x_{i},t_{i})|^{2}}}.$
These metrics represent the normalized differences between the model predictions and the ground truth values, providing a quantitative evaluation of model performance.
3 Numerical Experiments
In this section, we examine the proposed PINN scheme on a diverse set of prototypical stiff time-dependent PDEs with periodic boundary conditions, aiming to demonstrate its robustness and flexibility in handling a broad class of challenging equations. These benchmarks exhibit a wide range of physical and mathematical characteristics: from stiff dynamics (e.g., Kuramoto–Sivashinsky), to multi-species reaction–diffusion systems (e.g., Gray–Scott and Belousov–Zhabotinsky), and nonlinear dispersive waves (e.g., nonlinear Schrödinger equation). Such equations often involve strong nonlinearities, stiffness, and parameter sensitivities, posing significant challenges for data-driven methods such as PINNs. Nonetheless, our results illustrate the capacity of the proposed framework to overcome these difficulties and accurately approximate the solutions across these varied regimes. Detailed experimental results are presented in Section 3.1.
To begin, we employ Latin Hypercube Sampling to generate collocation points across the full spatio-temporal domain, including its boundary. To facilitate training and better incorporate the initial condition into the solution structure, we adopt the following transformation:
$\tilde{u}(t,\mathbf{x})=\exp(-Ct)*u_{0}(\mathbf{x})+t*u_{nn}(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\bar{\Omega};\quad\psi=\exp(-Ct)*u_{0}(\mathbf{x}),\,\phi=t.$
where $u_{0}(\mathbf{x})$ denotes the initial condition, and we test two values of $C\in\{0.1,1\}$, to assess how the decay rate affects learned dynamics. The transformed PINN $u_{nn}$ is trained to minimize a loss function that accounts for the PDE residuals, which now includes IC/BC. To rigorously assess the accuracy of our results, we use Chebfun, a spectral-style computing system in MATLAB that enables highly accurate function approximations. This allows us to make precise comparisons between the ground-truth solutions and the PINN-generated solutions.
For each benchmark, we not only present the main numerical results but also carry out detailed ablation studies to evaluate the contribution of each proposed technique. In each ablation experiment, we deactivate one component of the methodology at a time, keeping the remaining elements active under the same hyper-parameter configurations, to measure the relative $L^{2}$ error and relative $L^{1}$ error. The neural network architecture used throughout consists of 7 hidden layers with 32 neurons per layer and tanh activation functions. The initial weights are fixed across all ablation runs for consistency. Model training proceeds in two stages: first using the Adam optimizer kingma2014adam with a learning rate of $10^{-3}$ for $50,000$ steps, followed by fine-tuning via the L-BFGS-B optimizer with a learning rate of $1$ for $5,000$ steps. Each experiment uses $16,384$ interior collocation points, along with $128$ initial and $128$ boundary data points sampled via LHS. For PDE systems (e.g., Gray–Scott), we further apply a mini-batching strategy with $5$ batches and a batch size of $3,280$ to enhance training stability.
Overall, these extensive experiments highlight the effectiveness and adaptability of our PINN scheme in solving a wide variety of stiff time-dependent PDEs. The results provide not only strong empirical validation but also a reliable and reproducible benchmark for future developments in this area. For full experimental details and visualizations, see Section 3.1.
3.1 Allen-Cahn Equation $I$
The Allen-Cahn type PDE is a classical phase-field model that has been widely used to study phase separation phenomena bazant2017thermodynamic . It has numerous practical applications across a range of fields, including materials science allen1979microscopic ; shen2010numerical , biological systems hyman2014liquid ; takatori2015towards , and electrochemical systems horstmann2013precipitation ; tian2015electrochemically . We begin with the one-dimensional Allen-Cahn equation with periodic boundary conditions, a representative example that presents challenges for traditional PINN models, formulated as follows:
$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}u^{3}-\gamma_{2}u=0,\quad(t,x)\in(0,T)\times(a,b),$ (3)
$u(0,x)=u_{0}(x),\quad x\in[a,b],$
$u(t,a)=u(t,b),\quad t\in[0,T],$
$u_{x}(t,a)=u_{x}(t,b),\quad t\in[0,T],$
where $\gamma_{1},\gamma_{2}>0,T>0,a<b$ are prescribed constants. As $\gamma_{2}$ increases, the transition interface of the solutions is sharper, which makes it harder to solve the AC equation numerically. We first demonstrate the effectiveness of our scheme by testing on Allen-Cahn equation with the following initial condition and parameters, $u_{0}(x)=x^{2}\cos(\pi x)$, $T=1$, $a=-1$, $b=1$, $\gamma_{1}=0.001$, and $\gamma_{2}=5$.
Figure 1: Exact solution of the 1D Allen-Cahn Equation (Case I) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.18\times 10^{-3}$ and the relative $L^{1}$ error is $5.34\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 2: Best solutions of the Allen-Cahn Equation (Case I) $\gamma_{2}=5$. The resulting relative $L^{2}$ error is $1.18\times 10^{-3}$ and the relative $L^{1}$ error is $5.34\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
For this example, we first test the baseline PINN approach raissi2019physics . As shown in Table 1, the baseline PINN alone fails to accurately solve the Allen-Cahn equation, motivating the development of a more effective training pipeline. To address this, we employ all proposed techniques (except mini-batching) and conduct an ablation study by disabling individual components one at a time. Additionally, we introduce two different decay functions for the initial condition, $e^{-1t}$ and $e^{-0.1t}$, to investigate their influence on the solution. The results are summarized in Table 1. The full scheme, including self-adaptive loss balancing, achieves the best performance, reaching a relative $L^{2}$ error of $1.18\times 10^{-3}$ and relative $L^{1}$ error of $5.34\times 10^{-4}$. In particular, using the faster-decaying initial condition ($IC1$) leads to better results than the slower-decaying case ($IC2$), suggesting that the initial condition plays a more critical role during the early stages of training but becomes less dominant over time. The ablation results further highlight the importance of each methodological component. Removing any single part generally leads to performance degradation, with the most severe impact observed when the initial condition is omitted, resulting in a relative error close to $0.99$ in both $L^{1}$ and $L^{2}$ norms. Figure 1 illustrates the exact solution, the network prediction, and the absolute error distribution. As seen, the predicted solution matches the ground truth remarkably well, with errors concentrated at negligible levels. Figure 2 presents cross-sectional comparisons at different times, further confirming the excellent agreement between the predicted and true solutions across the entire time horizon. These results provide strong evidence for the effectiveness of the proposed training strategy and underscore the critical role of each pipeline component.
Table 1: Allen-Cahn Equation (Casse $I$ and Case $II$): Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance - Case $I$ Performance - Case $II$
PBC IC1 IC2 SA Rel. $L^{2}$ error Rel. $L^{1}$ error Rel. $L^{2}$ error Rel. $L^{1}$ error
$1.18\times 10^{-3}$ $5.34\times 10^{-4}$ $3.53\times 10^{-4}$ $2.14\times 10^{-4}$
$7.64\times 10^{-2}$ $1.66\times 10^{-2}$ $7.04\times 10^{-4}$ $3.26\times 10^{-4}$
$1.24\times 10^{-3}$ $6.10\times 10^{-4}$ $8.32\times 10^{-4}$ $3.72\times 10^{-4}$
$2.86\times 10^{-2}$ $8.20\times 10^{-3}$ $9.80\times 10^{-4}$ $3.91\times 10^{-4}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$1.05\times 10^{-2}$ $2.65\times 10^{-3}$ $5.11\times 10^{-2}$ $1.21\times 10^{-2}$
$1.86\times 10^{-3}$ $7.31\times 10^{-4}$ $5.03\times 10^{-2}$ $8.27\times 10^{-3}$
$1.88\times 10^{-2}$ $4.71\times 10^{-3}$ $4.58\times 10^{-2}$ $1.10\times 10^{-2}$
$4.34\times 10^{-3}$ $1.24\times 10^{-3}$ $3.03\times 10^{-2}$ $4.25\times 10^{-3}$
$5.12\times 10^{-1}$ $3.18\times 10^{-1}$ $4.02\times 10^{-1}$ $1.42\times 10^{-1}$
$9.99\times 10^{-1}$ $9.98\times 10^{-1}$ $5.11\times 10^{-1}$ $3.28\times 10^{-1}$
3.2 Allen-Chan, Case $II$
For our second example, we investigate the performance of PINNs in scenarios involving sharper moving interfaces. Specifically, we consider a modified Allen–Cahn equation with more oscillatory initial data:
$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}u^{3}-\gamma_{2}u=0,\quad(t,x)\in(0,1)\times(-1,1),$ (4)
$u(0,x)=x^{2}\sin(2\pi x),\quad x\in[-1,1],$
$u(t,-1)=u(t,1),\quad t\in[0,1],$
$u_{x}(t,-1)=u_{x}(t,1),\quad t\in[0,1],$
where $\gamma_{1}=0.001,\gamma_{2}=4$.
We first train the baseline PINN model, which produces a large relative $L^{2}$ error of $0.511$ and relative $L^{1}$ error of $0.328$, demonstrating its inadequacy in resolving the sharp interface dynamics (Table 1). However, once the initial condition enforcement and periodic boundary neural networks are incorporated, the performance improves drastically, achieving a relative $L^{2}$ error of $8.32\times 10^{-4}$ and relative $L^{1}$ error of $3.72\times 10^{-4}$. Further gains are realized when using the self-adaptive loss weighting strategy, although the improvement is marginal compared to the dominant contribution from enforcing the initial and boundary conditions. Interestingly, unlike the previous case, there is minimal performance difference between the schemes using $IC1$ and $IC2$, suggesting that the impact of initial condition decay becomes less significant when strong oscillations are already present at the beginning.
Figure 3: Exact solution of the Allen-Cahn Equation (Case II) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 4: Best solutions of the Allen-Cahn Equation (Case II) $\gamma_{2}=4$. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
3.3 Cahn-Hilliard Equation
Figure 5: Exact solution of the 1D Cahn-Hilliard Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $9.75\times 10^{-4}$ and the relative $L^{1}$ error is $5.65\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 6: Best solutions of the 1D Cahn-Hilliard Equation. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
The Cahn-Hilliard equation, a variant of the Allen-Cahn equation, also models phase separation. This process describes how the components of a binary fluid spontaneously separate into distinct domains, each composed of a single component cahn1958free ; kim2016basic ; lee2014physical ; miranville2017cahn . The equation is expressed as:
$u_{t}=\epsilon_{1}(-u_{xx}-\epsilon_{2}u_{xxxx}+(u^{3})_{xx}),\quad(t,x)\in(0,T]\times(-L,L),$ (5)
$u(0,x)=u_{0}(x),\quad x\in[-L,L],$
$u(t,-L)=u(t,L),\quad t\in[0,T],$
where $\epsilon_{1}=10^{-2},\epsilon_{2}=10^{-4},T=1,L=1$. We specify the initial condition as $u_{0}(x)=-\cos(2\pi x)$. This PDE involves higher-order derivatives, making it more challenging to solve compared to the Allen-Cahn equation.
For the ablation study in Section 2, the neural network is trained using the same hyperparameters outlined in Section 3. In this experiment, the full scheme yields the best results. Additionally, methods using initial condition IC2 outperform those with IC1. Except for the full schemes and the experiment with periodic neural networks for boundary conditions and initial condition IC2, none of the other approaches produce accurate solutions. The errors for all experiments are summarized in Table 2. The best results are illustrated in Figures 5 and 6.
Table 2: Cahn-Hilliard Equation: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance
PBC IC1 (C=1) IC2 (C=0.1) SA Rel. $L^{2}$ error Rel. $L^{1}$ error
$1.06\times 10^{-3}$ $5.61\times 10^{-4}$
$9.75\times 10^{-4}$ $5.65\times 10^{-4}$
$2.50\times 10^{-1}$ $5.87\times 10^{-2}$
$3.87\times 10^{-3}$ $1.91\times 10^{-3}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$3.48\times 10^{-1}$ $2.94\times 10^{-1}$
$2.37\times 10^{-1}$ $1.09\times 10^{-1}$
$4.07\times 10^{-1}$ $3.45\times 10^{-1}$
$1.55\times 10^{-2}$ $8.20\times 10^{-3}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.98\times 10^{-1}$
3.4 Kuramoto-Sivashinsky Equation
Kuramoto–Sivashinsky equation is a fourth-order nonlinear partial differential equation, known for its chaotic behavior kudryashov1990exact ; michelson1986steady . It has served as a pivotal model in the study of complex dynamical phenomena observed in various physical systems. Consequently, the Kuramoto–Sivashinsky equation provides an excellent case study to demonstrate how PINNs can effectively simulate chaotic dynamics. The equation is given by:
$u_{t}=-u_{xx}-u_{xxxx}-uu_{x},\quad(t,x)\in(0,T]\times(a,b),$ (6)
$u(0,x)=u_{0}(x),\quad x\in[a,b],$
$u(t,a)=u(t,b),\quad t\in[0,T],$
where $T=20,a=0,b=32\pi$, and $u_{0}(x)=\cos\frac{x}{16}(1+\sin\frac{x-1}{16})$. The parameters are the same as those mentioned in Section 3. The best results are shown in Figure 7 and 8:
Figure 7: Exact solution of the 1D Kuramoto-Sivashinsky Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.46\times 10^{-3}$ and the relative $L^{1}$ error is $1.02\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 8: Best solutions of the 1D Kuramoto-Sivashinsky Equation. The resulting relative $L^{2}$ error is $1.46\times 10^{-3}$ and the relative $L^{1}$ error is $1.02\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Table 3: Kuramoto-Sivashinsky Equation: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance
PBC IC1 (C=1) IC2 (C=0.1) SA Rel. $L^{2}$ error Rel. $L^{1}$ error
$4.04\times 10^{-2}$ $1.46\times 10^{-2}$
$1.46\times 10^{-3}$ $1.02\times 10^{-3}$
$1.68\times 10^{-2}$ $8.94\times 10^{-3}$
$1.30\times 10^{-2}$ $7.79\times 10^{-3}$
$9.84\times 10^{-3}$ $6.36\times 10^{-3}$
$2.50\times 10^{-3}$ $1.81\times 10^{-3}$
$2.13\times 10^{-1}$ $9.74\times 10^{-2}$
$2.00\times 10^{-1}$ $1.04\times 10^{-1}$
$1.54\times 10^{-2}$ $1.21\times 10^{-2}$
$9.60\times 10^{-3}$ $8.75\times 10^{-3}$
$2.11\times 10^{-1}$ $1.06\times 10^{-1}$
$3.01\times 10^{-1}$ $1.97\times 10^{-1}$
As in Table 3, solving this PDE system using a baseline PINN alone proves to be exceptionally challenging. The incorporation of periodic neural networks for boundary conditions, paired with initial condition IC2, consistently yields superior results compared to IC1. This suggests that, in contrast to the Allen-Cahn equation discussed in Section 3.1, the choice of initial condition plays a more significant role in the training process for this more complex system. This observation underscores the importance of carefully selecting both the boundary conditions and initial conditions when training neural networks for chaotic or higher-order PDE systems, where these factors contribute notably to the accuracy and stability of the solution.
3.5 Gray-Scott Equation
Reaction and diffusion of chemical species can produce a variety of patterns de1991turing ; nishiura1999skeleton , reminiscent of those often seen in nature. The Gray-Scott type system is one of classical mathematical models for chemical reactions gray1983autocatalytic ; gray1984autocatalytic ; liang2022reversible . The general irreversible Gray-Scott equations describe such reactions:
$U+2V$ $\longrightarrow 3V,$ (7)
$V$ $\longrightarrow P.$
This system is defined by two equations that describe the dynamics of two reacting substances:
$u_{t}=\epsilon_{1}u_{xx}+b(1-u)-uv^{2},\quad(t,x)\in(0,T]\times(-L,L),$ (8)
$v_{t}=\epsilon_{2}v_{xx}-(b+k)v+uv^{2},$
$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad\forall x\in[-L,L],$
$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad\forall t\in[0,T],$
where $T=20,L=50$, $\epsilon_{1}=1,\epsilon_{2}=0.01$ are diffusion rates, $b=0.02$ is the "feeding rate" that adds $U$, $k=0.0562$ is the "killing rate" that removes $V$. We set our initial conditions as:
$u_{0}(x)=1-\frac{\sin(\pi(x-50)/100)^{4}}{2},\quad v_{0}(x)=\frac{\sin(\pi(x-50)/100)^{4}}{4}.$
In this ablation study, we not only evaluate the performance of self-adaptive neural networks but also test our scheme with mini-batching on the Gray-Scott equations, which model the densities of two interacting species. For mini-batching, we use 5 batches, each containing 3,280 samples. As shown in Table 4, the results confirm that each proposed component improves the overall model performance. Omitting any one of these components leads to higher error rates. Notably, the errors for the $u-$equation are typically smaller than those for the $v-$equation, which aligns with our expectations. This is because the $v-$equation is sharper and more challenging to train compared to the $u-$equation. Additionally, experiments using initial condition IC2 consistently outperform those using IC1. Specifically, the full scheme with IC2 achieves the best results, with a relative $L^{2}$ error of $2.05\times 10^{-4}$ for $u-$equation and $8.46\times 10^{-4}$ for $v-$equation. These results suggest that IC2 is more effective for larger time-scale PDEs, as IC1 causes the initial condition to decay too quickly, reducing its influence during training. The predicted solutions from our best model are visualized in Figures 9 and 10, respectively.
Figure 9: Best solutions of the 1D Gray-Scott Equation (u species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 10: Best solutions of the 1D Gray-Scott Equation (v species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 11 and Figure 12 show the comparison of the best prediction against the reference solution. Note that the majority of this error is attributed to the sharpness of the PDEs.
Figure 11: Exact solution of the 1D Gray-Scott Eqaution (v specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 12: Exact solution of the 1D Gray-Scott Eqaution (v specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Table 4: Gray-Scott Equation: Relative $L^{2}$ errors of $u$ and $v$ equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance
PBC IC1 IC2 SA mini-batch Rel. $L^{2}$ error (u - eq.) Rel. $L^{2}$ error (v - eq.)
$8.98\times 10^{-4}$ $7.92\times 10^{-3}$
$2.05\times 10^{-4}$ $8.46\times 10^{-4}$
$4.93\times 10^{-3}$ $3.03\times 10^{-2}$
$2.86\times 10^{-3}$ $8.20\times 10^{-3}$
$9.01\times 10^{-3}$ $4.27\times 10^{-3}$
$7.59\times 10^{-3}$ $3.98\times 10^{-3}$
$5.11\times 10^{-3}$ $2.58\times 10^{-2}$
$1.23\times 10^{-3}$ $9.74\times 10^{-3}$
$8.09\times 10^{-1}$ $9.99\times 10^{-1}$
$2.06\times 10^{-1}$ $7.43\times 10^{-1}$
$5.37\times 10^{-1}$ $2.58\times 10^{-1}$
$1.15\times 10^{-1}$ $2.62\times 10^{-1}$
$9.53\times 10^{-2}$ $8.65\times 10^{-2}$
$9.21\times 10^{-2}$ $2.98\times 10^{-1}$
3.6 Belousov-Zhabotinsky Equation
A Belousov–Zhabotinsky reaction is one of a class of reactions that serve as a classical example of non-equilibrium thermodynamics. It typically involves the oxidation of organic compounds by bromine in an acidic medium zhabotinsky2007belousov ; cassani2021belousov . The system of PDEs takes the following form:
$u_{t}=\epsilon_{1}u_{xx}+u+v-uv-u^{2},$ (9)
$v_{t}=\epsilon_{2}v_{xx}+w-v-uv,$
$w_{t}=\epsilon_{1}w_{xx}+u-w,\quad(t,x)\in(0,T]\times[-L,L]$
$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad w(0,x)=w_{0}(x),\quad\forall x\in[-L,L]$
$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad w(t,-L)=w(t,L),\quad\forall t\in[0,T],$
where $T=2,L=1$, $\epsilon_{1}=10^{-5},\epsilon_{2}=2\times 10^{-5}$ are diffusion rates. We let the initial conditions be:
$u_{0}(x)=\exp(-100(x+0.5)^{2}),\quad v_{0}(x)=\exp(-100x^{2}),\quad w_{0}(x)=\exp(-100(x-0.5)^{2}).$
The system of PDEs involving three species provides a valuable opportunity to evaluate our proposed schemes using the techniques described in Section 2 for simulating multispecies dynamics. The results, summarized in Table 5, show that the full scheme consistently achieves the best overall performance, in line with previous experiments. Interestingly, the errors observed for methods using IC1 and those using IC2 are not significantly different. Both IC1 and IC2, when used independently, are capable of producing satisfactory results. However, when periodic neural networks are used to enforce boundary conditions, the performance declines and becomes less stable, even though self-adaptive training and mini-batching provide some improvements. This observation suggests that incorporating boundary conditions introduces additional complexity into the training process. Further investigation into their effects would be a promising direction for future research.
Figure 13: Best solutions of the 1D Belousov-Zhabotinsky equation (u specie). The resulting relative $L^{2}$ error is $3.49\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 14: Best solutions of the 1D Belousov-Zhabotinsky equation (v specie). The resulting relative $L^{2}$ error is $2.45\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 15: Best solutions of the 1D Belousov-Zhabotinsky equation (w specie). The resulting relative $L^{2}$ error is $2.50\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
The comparisons between the exact solutions of $(u,v,w)$ equations with corresponding best neural network predictions on the whole $(x,t)$ domain are shown in Figures 16, 17 and 18. Although there is a slight misalignment between each two of them, the model predictions closely match the corresponding numerical estimations.
Figure 16: Exact solution of the 1D Belousov-Zhabotinsky equation (u specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $3.49\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 17: Exact solution of the 1D Belousov-Zhabotinsky equation (v specie) Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.45\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 18: Exact solution of the 1D Belousov-Zhabotinsky equation (w specie) Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.50\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Table 5: Belousov-Zhabotinsky Equation: Relative $L^{2}$ errors of $u$, $v$ and $w$ equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance: relative $L^{2}$ error
PBC IC1 IC2 SA mini-batch u-equation v-equation w-equation
$3.49\times 10^{-3}$ $2.45\times 10^{-3}$ $2.50\times 10^{-3}$
$6.06\times 10^{-3}$ $4.47\times 10^{-3}$ $4.30\times 10^{-3}$
$7.08\times 10^{-3}$ $4.77\times 10^{-3}$ $4.24\times 10^{-3}$
$5.03\times 10^{-3}$ $3.10\times 10^{-3}$ $3.71\times 10^{-3}$
$9.17\times 10^{-3}$ $9.29\times 10^{-3}$ $8.89\times 10^{-3}$
$1.05\times 10^{-1}$ $1.78\times 10^{-2}$ $1.29\times 10^{-2}$
$2.22\times 10^{-2}$ $1.33\times 10^{-2}$ $1.52\times 10^{-2}$
$1.54\times 10^{-2}$ $9.17\times 10^{-3}$ $1.07\times 10^{-2}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$7.99\times 10^{-3}$ $5.05\times 10^{-3}$ $5.16\times 10^{-3}$
$9.39\times 10^{-3}$ $9.41\times 10^{-3}$ $9.85\times 10^{-3}$
$3.98\times 10^{-2}$ $2.15\times 10^{-2}$ $2.79\times 10^{-2}$
$4.48\times 10^{-2}$ $2.71\times 10^{-2}$ $3.22\times 10^{-2}$
$7.87\times 10^{-2}$ $3.95\times 10^{-2}$ $5.48\times 10^{-2}$
3.7 Nonlinear Schröedinger Equation
In theoretical physics, nonlinear Schröedinger equation is a nonlinear PDE, applicable to classical and quantum mechanics kato1987nonlinear ; kevrekidis2001discrete ; fibich2015nonlinear . The dimensionless equation of the classical field is
$u_{t}=iu_{xx}+i|u|^{2}u,\quad(t,x)\in[0,2]\times[-\pi,\pi],$ (10)
$u(0,x)=u_{0}(x),\quad x\in[-\pi,\pi],$
$u(t,-\pi)=u(t,\pi),\quad t\in[0,2],$
where we let
$u_{0}(x)=\frac{2}{2-\sqrt{2}\cos(x)}-1.$
Using the same parameter settings and training steps with prior experiments, we obtain the following results:
Figure 19: Exact solution of the real part of 1D Nonlinear Schroedinger Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.80\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 20: Best real-part solutions of the 1D Nonlinear Schroedinger Equation. The resulting relative $L^{2}$ error is $1.80\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 21: Exact solution of the imaginary part of 1D Nonlinear Schroedinger Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.58\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 22: Best imaginary-part solutions of the 1D Nonlinear Schroedinger Equation. The resulting relative $L^{2}$ error is $2.58\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Table 6 presents the results of all experiments. The full scheme achieves the best performance, with a relative $L^{2}$ error of real part of $1.80\times 10^{-4}$ and the relative $L^{2}$ error of imaginary part of $2.58\times 10^{-4}$. There is no significant difference in the errors between methods using IC1 and those using IC2. Moreover, without the use of self-adaptive neural networks or mini-batching, accurately obtaining the solution becomes challenging, highlighting the importance of both methods in ensuring more stable training.
Table 6: Nonlinear Schroedinger Equation: Relative $L^{2}$ errors of real and imaginary equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance: relative $L^{2}$ error
PBC IC1 IC2 SA mini-batch Real part Imaginary part
$1.80\times 10^{-3}$ $2.58\times 10^{-3}$
$1.84\times 10^{-3}$ $3.27\times 10^{-3}$
$7.34\times 10^{-2}$ $1.23\times 10^{-1}$
$8.32\times 10^{-2}$ $1.39\times 10^{-1}$
$1.10\times 10^{-2}$ $2.10\times 10^{-2}$
$1.70\times 10^{-2}$ $3.27\times 10^{-2}$
$1.00\times 10^{-1}$ $1.83\times 10^{-1}$
$9.56\times 10^{-2}$ $1.56\times 10^{-1}$
$8.09\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $8.78\times 10^{-1}$
$5.04\times 10^{-2}$ $1.21\times 10^{-1}$
$5.44\times 10^{-2}$ $7.26\times 10^{-2}$
$6.60\times 10^{-2}$ $1.05\times 10^{-1}$
$3.63\times 10^{-1}$ $5.06\times 10^{-1}$
4 Conclusion
We provide an ablation study of hard-constraints and self-adaptive loss weights to study the solution dependency of time-dependent PDEs on its initial conditions. The transformation provided by hard-constraints significantly alleviates the training challenges typically associated with stiff time-dependent PDEs subject. Our study demonstrates effectiveness across seven distinct types of such PDEs, encompassing scalar-valued outputs, vector-valued outputs, and even complex-valued solutions. This approach not only enhances training efficiency but also improves the accuracy and stability of the learned solutions, particularly in the presence of stiffness, where conventional PINNs often struggle to converge.
Although we provide an ablation study of hard-constraints and self-adaptive loss weights, our list of possible training remedies for PINNs is not exhaustive. Moreover, our tests should also include elliptical PDEs with high oscillatory coefficients. Future work includes extending the hard-cosntraint framework to two- and three-dimensional PDEs, which pose additional challenges in both architecture design and computational cost. Ongoing research also focuses on identifying optimal transformations $\psi$ and $\phi$, which govern the embedding of IC and BC into the network structure, with the goal of further enhancing performance and generalizability.
Acknowledgements
BH and MZ thank Prof. Yannis Kevrekids (JHU) for the discussion of well-conditioning of PDE solution maps, and Prof. George Karniadakis (Brown) for his suggestion on training enhancement methods for PINNs. BH developed and implemented the algorithm, ran the simulation and analyzed the data. UB and MZ designed the reserach. All authors wrote the paper. MZ is supported by NSF-AoF grant #$\#2225507$ and Ralph E. Powe Junior Faculty Enhancement Award from ORAU for FY24242424.
Supplementary Material
The source code, initial training data, initial weights, Chebfun code, and ground truth solutions used for benchmarking will provided in https://github.com/baolihao/Training-PINNs-with-Hard-Constraints-and-Adaptive-Weights. The list of hyperparameters and computational resources used to produce the results will be discussed later in this appendix.

### v3

@@
-Training PINNs with Hard Constraints and Adaptive Weights: An Ablation Study
+Stability in Training PINNs for Stiff PDEs: Why Initial Conditions Matter
 Abstract
-Training Physics-Informed Neural Networks (PINNs) to solve stiff time-dependent partial differential equations (PDEs) remains a significant challenge.
-The main difficulties lie in precisely enforcing initial conditions and balancing the various loss components.
-In stiff PDEs, where solutions show rapid transitions or sharp gradients, this balance becomes especially difficult.
-The wide range of scales in the dynamics can cause certain losses to dominate, leading to unstable or inefficient training.
-This paper presents a comprehensive ablation study focused on two pivotal training schemes: the enforcement of hard constraints for initial and boundary conditions, and the implementation of adaptive weights.
-We specifically examine their impact on the performance of PINNs when applied to stiff time-dependent PDEs from materials science and mathematical biology applications.
-We conduct extensive numerical experiments across a diverse range of time-dependent PDEs from Allen-Cahn, Cahn-Hillard, to Gray-Scott systems.
-We further discuss the implications of our findings for improving the robustness and efficiency of PINN training, particularly in settings where accurate representation of initial conditions and balanced loss contributions are paramount.
+Training Physics-Informed Neural Networks (PINNs) on stiff time-dependent PDEs remains highly unstable.
+Through rigorous ablation studies, we identify a surprisingly critical factor: the enforcement of initial conditions.
+We present the first systematic ablation of two core strategies, hard initial-condition constraints and adaptive loss weighting.
+Across challenging benchmarks (sharp transitions, higher-order derivatives, coupled systems, and high frequency modes), we find that exact enforcement of initial conditions (ICs) is not optional but essential.
+Our study demonstrates that stability and efficiency in PINN training fundamentally depend on ICs, paving the way toward more reliable PINN solvers in stiff regimes.
 1 Introduction
 Time-dependent partial differential equations (PDEs) constitute the mathematical foundation of numerous scientific and engineering disciplines, serving as essential tools for modeling a diverse range of physical phenomena.
-Their applications span from the intricate dynamics of fluid flow, as described by the Navier–Stokes equations, to the complex interplay of chemical reactions in biological systems, captured by reaction–diffusion models.
-Time-dependent PDEs offer a powerful framework for understanding and predicting the evolution of systems over time whitham2011linear ; burgers1948mathematical ; allen1979microscopic ; shen2010numerical ; bazant2017thermodynamic ; horstmann2013precipitation ; tian2015electrochemically ; hyman2014liquid ; takatori2015towards ; cahn1958free ; lee2014physical ; miranville2017cahn ; kudryashov1990exact ; michelson1986steady ; de1991turing ; nishiura1999skeleton ; zhabotinsky2007belousov ; cassani2021belousov ; kato1987nonlinear ; kevrekidis2001discrete ; fibich2015nonlinear .
-In particular, the imposition of periodic boundary conditions allows for the simulation of phenomena over effectively infinite physical domains, providing valuable insights into the intrinsic behavior of physical systems.
-However, even the numerical solution of these PDEs, especially in scenarios involving stiffness or complex non-linearities, can be exceedingly challenging, often demanding significant computational resources and specialized numerical techniques.
-The advent of machine learning has ushered in a new era of scientific computing, offering powerful tools for data-driven modeling and analysis.
-Machine learning methodologies have shown remarkable success in various science/engineering fields, revolutionizing areas such as protein folding prediction PF2019 , drug discovery and development CARRACEDOREBOREDO20214538 , and even optimizing fundamental algorithmic processes such as matrix-vector multiplication MMML2022 .
-However, traditional machine learning approaches, which rely on vast datasets, can prove impractical in scientific contexts where data acquisition is expensive, time consuming, or simply infeasible.
-Furthermore, purely data-driven models often lack physical fidelity and model interpretability, as they may fail to capture the underlying physical principles governing the systems they aim to describe.
-Physics-informed machine learning (PIML), specifically physics-informed neural networks (PINNs), address these limitations by seamlessly integrating physical laws into the learning process, significantly reducing the amount of training data needed.
-By embedding constraints derived from the governing PDEs directly into the neural network architecture, PINNs enable accurate predictions even with sparse and noisy data yang2019physics ; chen2021solving ; raissi2018hidden ; raissi2019physics ; liu2019multi ; liu2021dual ; jin2021nsfnets ; cai2021deepm ; rad2020theory ; zhu2021machine ; chen2022using ; wang2023pinnsbased .
-Despite the promising applications, effective training of PINNs, especially for stiff time-dependent PDEs, remains a significant challenge.
-The inherent difficulty in propagating information from initial and boundary data points into the computational domain, coupled with the need to balance multiple loss terms, often leads to convergence issues and suboptimal solutions wang2020understanding ; wang2020and ; mcclenny2020self ; coutinho2022physics .
-This paper presents an ablation study of training algorithms for PINNs applied to stiff partial differential equations (PDEs), focusing specifically on two schemes: hard constraints and self-adaptive loss weights mcclenny2020self .
+Their applications span from the intricate dynamics of fluid flow, as described by the Navier–Stokes equations, to the complex interplay of chemical reactions in biological systems, captured by reaction–diffusion models (Whitham, 2011; Burgers, 1948; Allen & Cahn, 1979; Shen & Yang, 2010; Bazant, 2017; Horstmann et al., 2013; Tian et al., 2015; Hyman et al., 2014; Takatori & Brady, 2015; Cahn & Hilliard, 1958; Lee et al., 2014; Miranville, 2017; Kudryashov, 1990; Michelson, 1986; De Kepper et al., 1991; Nishiura & Ueyama, 1999; Zhabotinsky, 2007; Cassani et al., 2021; Kato, 1987; Kevrekidis et al., 2001; Fibich, 2015).
+However, solution of stiff time-dependent PDEs can be exceedingly challenging to obtain and often requires substantial computational resources and specialized numerical techniques.
+Machine learning methodologies have demonstrated remarkable success across various scientific and engineering fields, transforming areas such as protein folding prediction (Jumper et al., 2021), drug discovery and development (Carracedo-Reboredo et al., 2021), and even the optimization of fundamental algorithmic processes such as matrix-vector multiplication (Fawzi et al., 2022).
+However, traditional machine learning approaches, which rely on large datasets, can be impractical in scientific contexts where data acquisition is expensive, time consuming, or simply infeasible.
+Moreover, purely data-driven models often lack physical fidelity and interpretability, as they may fail to capture the underlying physical principles governing the systems they aim to describe.
+Physics-informed machine learning (PIML), and in particular physics-informed neural networks (PINNs), address these limitations by integrating physical laws directly into the learning process.
+This significantly reduces the amount of training data required.
+By embedding constraints derived from governing PDEs into the neural network architecture, PINNs enable accurate predictions even when data are sparse or noisy (Yang et al., 2019; Chen et al., 2021; Raissi & Karniadakis, 2018; Raissi et al., 2019; Liu & Wang, 2019, 2021; Jin et al., 2021; Cai et al., 2021; Rad et al., 2020; Zhu et al., 2021; Chen et al., 2022; Wang et al., 2023).
+Despite their promise, effectively training PINNs for stiff time-dependent PDEs remains a major challenge.
+The difficulty of propagating information from initial and boundary conditions into the computational domain, combined with the need to balance multiple loss terms, often results in convergence issues and suboptimal solutions (Wang et al., 2020a, b; McClenny & Braga-Neto, 2020; Coutinho et al., 2023).
+This paper presents a complete study of investigating the importance of the training components for PINNs on solving stiff time-dependent PDES.
+First, we conduct an ablation study of training algorithms for PINNs applied to stiff time-dependent PDEs, focusing specifically on two schemes: hard constraints and self-adaptive loss weights (McClenny & Braga-Neto, 2020).
 The hard-constraint scheme directly embeds information about initial and boundary conditions into the neural network architecture.
-This transformation reduces the effects of spectral bias from training various kinds of loss terms, leading to improved solution accuracy wang2020and .
-PINN’s training can treated as a regularized regression, where initial and boundary conditions, along with observational data, serve as training data, and the physics-informed loss acts as regularization.
-However, the multi-objective nature of the loss function and the sensitivity of PDE solutions to initial conditions can significantly hinder the search for global minimizers.
-By explicitly incorporating boundary and initial condition data into the network transformation, we streamline the training process, particularly for stiff time-dependent PDEs, where accurately capturing the initial conditions is critical.
-In addition, we investigate the use of adaptive loss weights to alleviate the spectral bias from the multi-objective loss in PINN training mcclenny2020self .
-Through our ablation study of these training schemes, we arrive at the conclusion that hard constrain transformation provides better improvement.
-We emphasize that hard constraints and self-adaptive weights are not the only effective strategies for handling stiff PDEs.
-These methods can be integrated with other advanced PINN techniques, such as causal PINNs wang2020and , time-marching PINNs wight2020solving , residual-based-attention (RBA) PINNs ANAGNOSTOPOULOS2024116805 , curriculum training krishnapriyan2021characterizing , and co-training PINNs zhong2024label , to achieve synergistic improvements in prediction accuracy and robustness.
-Through a rigorous ablation study, we demonstrate the efficacy of the proposed approach across a range of stiff time-dependent PDEs, offering valuable insights into optimal training strategies for PINNs in these challenging scenarios.
+This transformation reduces the effects of spectral bias from training a multi-objective loss, leading to improved solution accuracy (Wang et al., 2020b).
+In addition, we investigate the use of adaptive loss weights to alleviate spectral bias in PINN training (McClenny & Braga-Neto, 2020).
+Through our ablation study of these training schemes, we conclude that the hard-constraint transformation provides the best improvement.
+Next, we provide a Neural Tangent Kernel perspective on how the hard-constrained transformation is reducing the spectral bias in PINN training.
+We emphasize that hard constraints is not the only effective strategies for handling stiff PDEs.
+Such method can be integrated with other advanced PINN techniques, such as causal PINNs (Wang et al., 2020b), time-marching PINNs (Wight & Zhao, 2020), residual-based-attention (RBA) PINNs (Anagnostopoulos et al., 2024), curriculum training (Krishnapriyan et al., 2021), and co-training PINNs (Zhong et al., 2024), to achieve synergistic improvements in prediction accuracy and robustness.
+The rest of the paper is structured as follows.
+We compare our methodology to other related works in section 1.1; in section 2, we discuss the hard-constrained framework and other training methodologies used in our ablation and comparison study; we presented a detailed discussion on the theoretical mechanism behind the hard-constrained transformation in section 3.1; we presented the results (in summary) from four major categories of examples in section 4; we conclude our paper in 5, pointing out several future directions for this line of research.
+Since the examples we conducted are too extensive, the additional details are presented in Appendix section B.
+Figure 1: 77 Benchmark Results: Truth vs HC-PINN vs Absolute Error.
 1.1 Related Work
@@
 However, challenges arise, particularly when dealing with “stiff” PDEs, where explicit numerical methods require extremely small time steps for stability.
-A primary concern is the imbalance between the fitting for initial data and residual components within the PINN loss function wight2020solving ; shin2020convergence ; wang2020understanding ; wang2020and .
-This imbalance often leads to a gradient descent that favors minimization of residual loss over data fitting loss, hindering convergence to accurate solutions.
+A primary concern is the imbalance between the fitting of initial data and residual components within the PINN loss function (Wight & Zhao, 2020; Shin et al., 2020; Wang et al., 2020a, b).
+This imbalance often leads gradient descent to prioritize minimizing the residual loss over data fitting loss, hindering convergence to accurate solutions.
 This issue is exacerbated in forward problems, where data is primarily concentrated at initial and boundary conditions, leaving limited information within the PDE domain.
-Specifically, in time-dependent problems, PINNs struggle to propagate information from initial conditions to subsequent time steps, a phenomenon widely documented in the literature wight2020solving ; mcclenny2020self ; krishnapriyan2021characterizing ; wang2022respecting ; haitsiukevich2022improved .
+Specifically, for time-dependent problems, PINNs struggle to propagate information from initial conditions to subsequent time steps, a phenomenon widely documented in the literature (Wight & Zhao, 2020; McClenny & Braga-Neto, 2020; Krishnapriyan et al., 2021; Wang et al., 2022; Haitsiukevich & Ilin, 2022).
 Several strategies have been proposed to address these challenges.
-One approach, introduced by Wight and Zhao wight2020solving , involves time-marching, where the time domain is divided into smaller intervals, and PINNs are sequentially trained on each interval, starting from the initial condition.
+One approach, introduced by Wight & Zhao (2020), involves time-marching, where the time domain is divided into smaller intervals, with PINNs being sequentially trained on each interval, starting from the initial condition.
 While effective, this method can be computationally expensive due to the need to train multiple networks.
-To improve information propagation, researchers have emphasized the importance of weighting initial condition data and residual points near the initial time wight2020solving ; mcclenny2020self .
-McClenny and Konstantinidis mcclenny2020self demonstrated that their self-adaptive PINNs can autonomously learn these weights during training, particularly in scenarios with complex initial conditions, such as the wave equation.
-Further advancements include the “sequence-to-sequence” approach proposed by Krishnapriyan et al.
-krishnapriyan2021characterizing , which sequentially learns each time step in a time-marching scheme.
-Wang et al.
-wang2022respecting identified the lack of “causality” in standard PINN training and proposed a corrective weighting scheme that gradually prioritizes data from early to later times, leveraging information from the PDE residue.
-This approach shares similarities with the self-adaptive weighting mechanism of McClenny and Konstantinidis mcclenny2020self .
-Additionally, Haitsiukevich et al.
-haitsiukevich2022improved introduced a pseudo-label approach, resembling self-training, to further enhance PINN convergence and accuracy.
-In zhong2024label , a mixed training scheme using labal-propagation between data points and collocation points is introduced and tested.
-For stiff parabolic PDEs such as viscous Burgers with small viscosity, the authors in coutinho2022physics proposed an data-driven artificial viscosity scheme to produce stable solutions and accurate to capture sharp transitions in the solutions.
+To improve information propagation, researchers have emphasized the importance of weighting initial condition data and residual points near the initial time (Wight & Zhao, 2020; McClenny & Braga-Neto, 2020).
+McClenny & Braga-Neto (2020) demonstrated that their self-adaptive PINNs can autonomously learn these weights during training, particularly in scenarios with complex initial conditions, such as the wave equation.
+Further advancements include the “sequence-to-sequence” approach (Krishnapriyan et al., 2021), causality PINN (Wang et al., 2022), artificial viscosity PINN (Coutinho et al., 2023), Co-training PINN (Zhong et al., 2024).
 Beyond loss weighting and time marching, structural modifications to PINNs have also been explored.
-Braga-Neto et al.
-braganeto2023characteristicsinformed introduced Characteristic-Informed Neural Networks (CINNs) for transport equations, where PDE information is integrated into the neural network architecture.
-A similar approach, albeit for Dirichlet boundary conditions, was proposed by Lagaris et al.
-Lagaris_1998 .
-Our method builds upon these architectural modifications, with a focus on enhancing stability and ensuring proper propagation of information, especially for stiff time-dependent PDEs.
-We carefully design our architecture to address the challenges of information propagation and loss balancing, aiming to improve the robustness and efficiency of PINN training.
+Braga-Neto (2022) introduced Characteristic-Informed Neural Networks (CINNs) for transport equations, where PDE information is integrated into the neural network architecture.
+A similar approach, albeit for Dirichlet boundary conditions, was proposed in (Lagaris et al., 1998).
+Our method builds upon these architectural modifications, with a focus on enforcing (near)-exact fit of the initial condition using hard-constrained transformation.
 2 Methodology
-We consider the following setup for a family of time dependent PDEs.
-Let $u$ be an unknown function defined on $[0,T]\times\Omega$ with $\Omega\subset\mathcal{R}^{d}$.
+We consider the following setup for general time dependent PDEs.
+Let $u$ be an unknown function defined on $0,T]\times\bar{\Omega}$ with $\Omega\subset\mathbb{R}^{d}$.
 Here the physical domain $\Omega$ comes with a Lipschitz boundary $\partial\Omega$ and $\bar{\Omega}=\Omega\cup\partial\Omega$.
-Let $\mathcal{P}$ be an operator which can include differentiation and integration.
-For example, for the $1D$ heat equation $u_{t}-\lambda u_{xx}=0$, $\mathcal{P}=\lambda\frac{\partial^{2}}{\partial x^{2}}$.
-Notice that $\mathcal{P}$ can be nonlinear.
-Let $\mathcal{B}$ be another operator on the boundary, that is, for Dirichlet boundary condition $u(t,x)=g(t,x)$ for $(t,x)\in[0,T]\times\partial\Omega$, $\mathcal{B}=\mathcal{I}$ (the identity).
-Then we say that $u$ is a solution of an PDE, if $u$ satisfies the following
-$\begin{cases}u_{t}-\mathcal{P}(u)(t,\mathbf{x})&=f(t,\mathbf{x}),\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ u(0,\mathbf{x})&=u_{0}(\mathbf{x}),\quad\mathbf{x}\in\bar{\Omega},\\ \mathcal{B}(u)(t,\mathbf{x})&=g(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega,\\ \end{cases}$ (1)
-Here the functions $u_{0}$, $f$, and $g$ are user-inputs and are assumed to satisfy desired regularity so that the existence and uniqueness of solutions for such a PDE is guaranteed.
-We also assume that the compactibility condition, that is, $g(0,\mathbf{x})=u_{0}(\mathbf{x})$ for $\mathbf{x}\in\partial\Omega$, is satisfied to prevent any ill-conditioning of the PDE problem.
+And we denote $\mathbf{x}=\begin{pmatrix}x_{1}&\cdots&x_{d}\end{pmatrix}^{\top}\in\Omega$.
+Then we say that $u$ is a solution of a time-dependent PDE, if $u$ satisfies the following
+$\begin{cases}u_{t}(t,\mathbf{x})+\mathcal{P}[u](t,\mathbf{x})&=f(t,\mathbf{x}),\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ u(0,\mathbf{x})&=u_{0}(\mathbf{x}),\quad\mathbf{x}\in\bar{\Omega},\\ \mathcal{B}[u](t,\mathbf{x})&=g(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega,\\ \end{cases}.$ (1)
+Here the operator $\mathcal{P}$ has order $K$ and it is defined as
+$\mathcal{P}[u]=h(u,\{\partial^{k}_{x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{d}^{j_{d}}}u\}_{k=1}^{K}),\quad\text{$h$ is any multivariate function},$
+where $\partial^{k}_{x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{d}^{j_{d}}}u$ is a set that contains all of partial derivatives of $u$ with respective to $\mathbf{x}$ of $k^{th}$ order in the multi variate sense, i.e., for the powers of each coordinate, we have
+$j_{1},j_{2},\cdots,j_{d}\geq 0\quad\text{and}\quad j_{1}+j_{2}+\cdots+j_{d}=k.$
+Furthermore $\mathcal{B}$ is an operator defined for $(t,\mathbf{x})$ on the boundary.
+We also assume that the functions $u_{0}$, $f$, and $g$ are user inputs that satisfy the desired regularity so that the existence and uniqueness of solutions for such a PDE are guaranteed.
+We also assume that the compatibility condition, $g(0,\mathbf{x})=u_{0}(\mathbf{x})$ for $\mathbf{x}\in\partial\Omega$, is satisfied to prevent any ill-conditioning of the solution.
 Types of Boundary Conditions: one can consider many kinds of boundary conditions as follows
-$\mathcal{B}(u)(t,\mathbf{x})$ $=u(t,\mathbf{x}),\quad\text{Dirichlet type},$
-$\mathcal{B}(u)(t,\mathbf{x})$ $=\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Neumann type, $\mathbf{n}$ is the outward normal vector to $\partial\Omega$},$
-$\mathcal{B}(u)(t,\mathbf{x};a,b)$ $=au(t,\mathbf{x})+b\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Robin type}.$
-The periodic boundary condition gives $d$ different equations, and they are given as follows,
-$\mathcal{B}(u)(t,\mathbf{x})=u(t,\mathbf{x})-u(t,\mathbf{x}+P\mathbf{e}_{i})=0,\quad i=1,\cdots,d.$ (2)
-Here $P$ is the period and $\mathbf{e}_{i}$ is the $i^{th}$ standard basis vector in $\mathcal{R}^{d}$, that is,
+$\mathcal{B}[u](t,\mathbf{x};a,b)=au(t,\mathbf{x})+b\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x}),\quad\text{Robin type},$
+where $\mathbf{n}$ is the outward normal vector to $\partial\Omega$ and $a,b$ are two fixed constants.
+Special values of $a$ and $b$ can give two other BC cases, such as when $b=0$ gives $\mathcal{B}[u](t,\mathbf{x})=u(t,\mathbf{x})$ (Dirichlet) and $a=0$ gives the $\mathcal{B}[u](t,\mathbf{x})=\frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x})$ (Neumann).
+The periodic boundary condition gives $d$ different equations, and they are,
+$\mathcal{B}[u](t,\mathbf{x})=u(t,\mathbf{x})-u(t,\mathbf{x}+P\mathbf{e}_{i})=0,\quad i=1,\cdots,d.$ (2)
+Here $P$ is the period and $\mathbf{e}_{i}$ is the $i^{th}$ standard basis vector in $\mathbb{R}^{d}$, that is,
 $\mathbf{e}_{i}=\underbrace{\begin{bmatrix}0&\cdots&0&1&0&\cdots&0\end{bmatrix}^{\top}}_{\text{$1$ is the in the $i^{th}$ position.}}.$
@@
 For example, we can have $\Omega=\Gamma_{1}\cup\cdots\cup\Gamma_{K}$ with each $\Gamma_{i}$ defines a different kind of BC.
-In this paper, we will mainly focus on the periodic type of boundary conditions, however other kinds of BCs can be
-Classical numerical techniques such as the Finite Difference Method (which approximates derivatives on a point mesh), Finite Element Method (utilizing basis approximation on a triangular mesh with weak formulation), and Spectral Method (employing Fourier basis with FFT on a point mesh) are widely utilized in scientific computations.
-However, recent advancements in scientific machine learning integrate the principles of physics, particularly the understanding of PDEs, into the training process of machine learning models.
-This integration enables the development of PINNs to solve for $u$ as follows: one tries to find an approximate solution from a set of deep neural networks (of the same depth, same number of neurons on each hidden layer, and same activation functions on each layer) $\mathcal{H}_{\text{NN}}$, which is a minimizer of the following loss functional
-$\text{Loss}(u_{nn})=\text{Data Loss}+\lambda*\text{Physical Loss}$
+In this paper, we will mainly focus on the periodic type of boundary conditions.
+Recent advancements in scientific machine learning integrate the principles of Physics, particularly the Physics-derived PDEs, into the training process of machine learning models.
+This integration enables the development of PINNs to solve for $u$ as follows: find an approximate solution from a set of deep neural networks $\mathcal{H}_{\text{NN}}$, where each network has the same depth, the same number of neurons on each hidden layer, and the same activation functions on each layer, that minimizes the following loss functional
+[…]
 where $\lambda$ is a regularization parameter, and the two losses are as follows
-Data $\text{Data Loss}=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
-$\quad+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2}$
-Physical $\text{Physical Loss}=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}$
+$\begin{cases}&\text{Data Loss}(u_{nn})=\frac{1}{N_{Data}}\sum_{i=1}^{N_{Data}}|u_{nn}(t^{Data},\mathbf{x}_{i}^{Data})-u(t^{Data},\mathbf{x}_{i}^{Data})|^{2}\\ &\text{IC Loss}(u_{nn})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2},\\ &\text{BC Loss}(u_{nn})=\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2},\\ &\quad\text{or}\\ &\text{BC Loss}(u_{nn})=\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}|(\mathcal{B}[u_{nn}]-g)(t_{j}^{BC},\mathbf{x}_{j}^{BC})|^{2},\\ &\text{PDE Loss}(u_{nn})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\frac{\partial u}{\partial t}+\mathcal{P}[u_{nn}]-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}\end{cases}$
 for $u_{nn}\in\mathcal{H}_{\text{NN}}$.
-Here $\{(t^{CL},\mathbf{x}^{CL})_{i}\}_{i=1}^{N_{CL}}\in(0,T]\times\Omega$ are called collocation points, $\{(0,\mathbf{x}^{IC})_{i}\}_{i=1}^{N_{IC}}\in\{0\}\times\bar{\Omega}$ are the initial condition points, and $\{(t^{BC},\mathbf{x}^{BC})_{i}\}_{i=1}^{N_{BC}}\in[0,T]\times\partial\Omega$ are the boundary condition points.
+Here $\{(t^{CL},\mathbf{x}^{CL})_{i}\}_{i=1}^{N_{CL}}\in(0,T]\times\Omega$ are called collocation points, $\{(0,\mathbf{x}^{IC})_{i}\}_{i=1}^{N_{IC}}\in\{0\}\times\bar{\Omega}$ are the initial condition points, $\{(t^{BC},\mathbf{x}^{BC})_{i}\}_{i=1}^{N_{BC}}\in[0,T]\times\partial\Omega$ are the boundary condition points, and $\{(t^{Data},\mathbf{x}^{Data})_{i}\}_{i=1}^{N_{Data}}\in(0,T]\times\Omega$ are known data points (they can be noisy).
 The minimizer, denoted as $u_{NN}$, will an be approximate solution to (1).
-2.1 Hard Constrains
-The main motivation for us to consider hard constrains transformation for PINNs is the possible ill-conditioning of the PDE solution map’s dependence on IC and BC.
-Traditional numerical methods do not need to encounter this kind of difficulty, as the IC data is exactly satisfied in their solutions.
-PINNs, on the other hand, use $L^{2}$-loss to fit the IC/BC and PDE residual data, which causes the solutions to be highly sensitive to the tightness of fit for the IC in the training (as pointed out in mcclenny2020self ).
+2.1 Hard Constraints
+The main motivation for us to consider hard constrains transformation for PINNs is the possible ill-conditioning of the PDE solution operator’s dependence on IC and BC.
+Traditional numerical methods do not encounter this kind of difficulty, as the IC data is exactly satisfied at the starting point of the solution loop.
+PINNs, on the other hand, use $L^{2}$-loss to fit the IC/BC and PDE residual data, which causes the solutions to be highly sensitive to the tightness of fit for the IC in the training, as pointed out in (McClenny & Braga-Neto, 2020).
 By transforming the need of fitting IC/BC conditions out of PINNs, it reduces the spectral bias in the training.
 Therefore, We consider the following transformation
-$\tilde{u}(t,\mathbf{x})=\psi(t,\mathbf{x})+\phi(t,\mathbf{x})u_{nn}(t,\mathbf{x})$
+$\tilde{u}(t,\mathbf{x})=\psi(t,\mathbf{x})+\phi(t,\mathbf{x})u_{nn}(t,\mathbf{x}),$ (3)
 Here the functions $\psi$ and $\phi$ are smooth functions with the following property
-$\psi(0,\mathbf{x})=u_{0}(\mathbf{x}),\quad\phi(0,\mathbf{x})=0,\quad\text{for $\mathbf{x}\in\bar{\Omega}$},$
-where in the case of periodic BC, we require $\psi$ and $\phi$ to be both $P$-periodic; for other BC types, we instead require $\phi$ and $\psi$ to satisfy
-$\psi(t,\mathbf{x})=g(t,\mathbf{x}),\quad\phi(t,\mathbf{x})=0,\quad\text{for $(t,\mathbf{x})\in[0,T]\times\partial\Omega$}.$
-When such transformation is used, then the training for $u_{nn}$ is changed, due to the fact that IC/BC can be built into the PDE loss, i.e.
-$\text{Loss}(u_{nn})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\tilde{u}_{t}-\mathcal{P}(\tilde{u}))(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2},\quad\text{for $\tilde{u}=\psi+\phi u_{nn}$}.$
-We have reduced the types of losses in the training; however the information about IC and BC has then moved into the PDE residual loss through $\psi$ and $\phi$.
-The difficulty in training the new PDE residual loss has been increased.
-Hence, other training enhancement techniques can be used to improve the new PDE residual training.
+$\begin{cases}&\psi(0,\mathbf{x})=u_{0}(\mathbf{x}),\quad\psi(0,\mathbf{x})=0,\quad\text{for $\mathbf{x}\in\bar{\Omega}$},\\ &\psi(t,\mathbf{x})=g(t,\mathbf{x}),\quad\psi(t,\mathbf{x})=0,\quad\text{for $(t,\mathbf{x})\in[0,T]\times\partial\Omega$}.\end{cases}$ (4)
+In the case of periodic BC, we simply require $\psi$ and $\phi$ to be both $P$-periodic. When such transformation is used, then the training for $u^{nn}$ is updated as
+$\text{Loss}_{HC}(u_{nn})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}|(\tilde{u}_{t}+\mathcal{P}[\tilde{u}]-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2},\quad\tilde{u}=\psi+\phi u_{nn}.$
+We have successfully reduced the multi-objective loss to a single-objective; however the information about IC and BC has been moved into the PDE residual loss through $\psi$ and $\phi$.
 Periodic Neural Networks: even though we require $\psi$ and $\phi$ to be periodic, we still need to build the information of periodicity into $u_{nn}$, hence we consider the following composition
-$u_{nn}(t,x)=f_{nn}(v(t,x)),\quad\text{$f_{nn}$ is a neural network, and $x\in\mathcal{R}$},$
+$u_{nn}(t,x)=f_{nn}(v(t,x)),\quad\text{where $f_{nn}$ is a neural network and $x\in\mathbb{R}$}.$
+Here $v(t,x)=\begin{bmatrix}t&\mathcal{F}_{m}(x)\end{bmatrix}^{\top}$ with
+$\mathcal{F}_{m}(x)=\begin{bmatrix}1&\cos\left(\frac{2\pi}{P}x\right)&\sin\left(\frac{2\pi}{P}x\right)&\cdots&\cos\left(\frac{2\pi}{P}mx\right)&\sin\left(\frac{2\pi}{P}mx\right)\end{bmatrix}$
+where $m$ is a positive integer hyperparameter controlling the number of Fourier modes.
+With this construction, each trigonometric feature is exactly PP–periodic, and therefore any linear combination (and hence $u_{nn}$) will inherit PP–periodicity (Dong & Ni, 2021).
+For $\mathbf{x}=\begin{bmatrix}x&y\end{bmatrix}$, we can use the following
+$v(t,x,y)=\begin{bmatrix}t&\mathcal{F}_{m}(x)&\mathcal{F}_{m}(y)\end{bmatrix}$;
+similarly for higher dimensional $\mathbf{x}=\begin{bmatrix}x_{1}&\cdots&x_{d}\end{bmatrix}$, we have
+$v(t,\mathbf{x})=\begin{bmatrix}t&\mathcal{F}_{m}(x_{1})&\cdots&\mathcal{F}_{m}(x_{d})\end{bmatrix}$.
+2.2 Other Training Enhancements
+We also consider mini-batching (to handle large sample size) and self-adaptive loss weights (McClenny & Braga-Neto, 2020), please see sections A.1 and A.2 for details.
+3 Theoretical Foundation for Hard Constraints
+We discuss the well-conditioning condition for the hard-constraint transformation and the Neural Tangent Kernel (NTK) formulation for the training decay rates in this section.
+3.1 Well Condition of the Transformation
+We start with the minimal regularity conditions on the pair $(\psi,\phi)$ for a well-defined transformation.
+Theorem 1 (Well Conditioning).
+The transformation is well defined when $\phi$ and $\psi$ is $C^{1}$ in time and have $K^{th}$ order partial derivatives w.r.t $\mathbf{x}$.
+Moreover, $\phi_{t}(0,\mathbf{x})\neq 0$.
+Then $u_{nn}$ will satisfy a new PDE
+$\begin{cases}&\phi\partial_{t}u_{nn}+\mathcal{P}[\psi+\phi u_{nn}]+\phi_{t}u_{nn}=f-\psi_{t},\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ &u_{nn}(0,\mathbf{x})=\frac{u_{t}(0,\mathbf{x})-\psi_{t}(0,\mathbf{x})}{\phi_{t}(0,\mathbf{x})},\quad\mathbf{x}\in\bar{\Omega}.\end{cases}$
+Remark 1.
+Hence our choice of setting $\psi(t,\mathbf{x})=e^{-Ct}u_{0}(\mathbf{x})$ and $\phi(t,\mathbf{x})=t$ satisfies the well-conditioning requirement.
+However, if we are interested in obtaining the solution at $T\rightarrow\infty$, using $\phi(t,\mathbf{x})=t$ will cause the learning to deteriorate, since $t\rightarrow\infty$.
+It is better to use $\phi(t,\mathbf{x})=1-e^{-Ct}$ with $\psi(t,\mathbf{x})=e^{-Ct}u_{0}(\mathbf{x})$, so that the influence of $\phi$ on $u_{nn}$ will settle down to a fixed number.
+Therefore we will obtain (when $T\rightarrow\infty$)
+$u_{nn}(\infty,\mathbf{x})=\mathcal{P}^{-1}[f](\infty,\mathbf{x})=u(\infty,\mathbf{x}).$
+In the case of $\mathcal{P}$ being linear and $\phi(t,\mathbf{x})=\phi(t)$, we end up with a much simpler PDE
+$\phi\partial_{t}u_{nn}+\phi\mathcal{P}[u_{nn}]+\phi^{\prime}u_{nn}=f-\psi_{t}-\mathcal{P}(\psi),$
+If $\phi(t)\neq 0$ when $t\neq 0$, then it can be simplified further to
+$\partial_{t}u_{nn}+\mathcal{P}[u_{nn}]+\frac{\phi^{\prime}}{\phi}u_{nn}=\frac{f-\psi_{t}-\mathcal{P}(\psi)}{\phi}$
+We can see that $u_{nn}$ satisfies a similar PDE to the original one with the additional reaction term $\frac{\phi^{\prime}}{\phi}u_{nn}$ and updated forcing $\frac{f-\psi_{t}-\mathcal{P}(\psi)}{\phi}$, which contains the information of $u_{0}$ and $g$.
+3.2 Why Initial Conditions Matter?
+The hard constraints transformation consider both the boundary condition and initial condition, however our emphasis is on the initial condition as the focus of our paper is about solving stiff time-dependent PDEs with PINNs.
+Stiff time-dependent PDEs have important features that exhibit over a widely varying time scaling, hence making even explicit time-integration methods from traditional numerical methods difficult to capture these features.
+Therefore by removing the initial condition through hard constraint transformation, we achieve the following superior performance
+  • We implicitly enforce the time direction into the training of PINNs, reducing the spectral bias from training with multi-objective loss.
+  • We mimic the behavior of traditional numerical methods, as they start out with either exact or approximated initial condition in their time integration.
+  • The hard constraint transformation works similarly like an implicit time integrator.
+3.3 How Hard Constraints Work?
+Theorem 2.
+Assume that $\mathcal{P}$ is linear and $\phi(t,\mathbf{x})=\phi(t)\neq 0$ for $t>0$, then for a one-hidden layer network, we have for a network $u(\mathbf{y};\theta)$ with $\bm{\theta}\in\mathbb{R}^{D}$ being the network parameters, the following relationship is satisfied
+$u(\mathbf{y};\bm{\theta})=\mathbf{h}^{\top}(\mathbf{y})\bm{\theta},\quad\mathbf{h}=\begin{bmatrix}h_{1}(\mathbf{y})&\cdots&\mathbf{h}_{D}(\mathbf{y})\end{bmatrix}^{\top}$,
+where $\mathbf{y}=\mathbf{z}=(t,\mathbf{x})\in(0,T]\times\Omega$ or $\mathbf{y}=\bm{\zeta}=(0,\mathbf{x}^{IC})\in\{0\}\times\bar{\Omega}$. Therefore the two losses become
+$\text{Loss}_{s}(u(\mathbf{y};\bm{\theta}))$ $=\frac{1}{2N}|\mathbf{A}\bm{\theta}-\mathbf{f}|_{\ell^{2}(\mathbb{R}^{N})}^{2}+\frac{1}{2M}|\mathbf{C}\bm{\theta}-\mathbf{u}_{0}|_{\ell^{2}(\mathbb{R}^{M})}^{2},\quad\text{(soft constrained)}$
+$\text{Loss}_{h}(u(\mathbf{z};\bm{\theta}))$ $=\frac{1}{2N}|\bm{\Lambda}_{\phi}\mathbf{A}\bm{\theta}+\bm{\Lambda}_{\phi_{t}}\mathbf{B}\bm{\theta}-\tilde{\mathbf{f}}|_{\ell^{2}(\mathbb{R}^{N})}^{2},\quad\text{(hard constrained)}$
+where
+$\mathbf{A}=\begin{bmatrix}\mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{1})]\\ \vdots\\ \mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{N})]\end{bmatrix},\quad\mathbf{B}=\begin{bmatrix}\mathbf{h}^{\top}(\mathbf{z}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\mathbf{z}_{N})\end{bmatrix},\quad\text{and}\quad\mathbf{C}=\begin{bmatrix}\mathbf{h}^{\top}(\bm{\zeta}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\bm{\zeta}_{M})\end{bmatrix}$,
+where $\mathbf{x}^{IC}_{i}\in\bar{\Omega}$; and for the diagonal matrices $\bm{\Lambda}_{\phi}$ and $\bm{\Lambda}_{\phi_{t}}$
+$\bm{\Lambda}_{\phi}=\text{diag}(\phi(\mathbf{z}_{1}),\cdots,\phi(\mathbf{z}_{N}))\quad\text{and}\quad\bm{\Lambda}_{\phi_{t}}=\text{diag}(\phi_{t}(\mathbf{z}_{1}),\cdots,\phi_{t}(\mathbf{z}_{N})),$
+and the vectors
+$\mathbf{f}=\begin{bmatrix}f(\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})\end{bmatrix},\quad\tilde{\mathbf{f}}=\begin{bmatrix}f(\mathbf{z}_{1})-\mathcal{L}_{s}[\psi](\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})-\mathcal{L}_{s}[\psi](\mathbf{z}_{N})\end{bmatrix},\quad\text{and}\quad\mathbf{u}_{0}=\begin{bmatrix}u_{0}(\mathbf{x}_{1}^{IC})\\ \vdots\\ u_{0}(\mathbf{x}_{M}^{IC})\end{bmatrix}$.
+Therefore the Hessians of the losses are
+$H_{s}(\bm{\theta})=\frac{1}{N}\mathbf{A}^{\top}\mathbf{A}+\frac{1}{M}\mathbf{C}^{\top}\mathbf{C},\quad H_{h}(\bm{\theta})=\frac{1}{N}(\mathbf{A}^{\top}\bm{\Lambda}_{\phi}^{2}\mathbf{A}+2\mathbf{A}^{\top}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\phi_{t}}\mathbf{B}+\mathbf{B}^{\top}\bm{\Lambda}_{\phi_{t}}^{2}\mathbf{B}),$
+Then the Hessians are bounded as
+$|H_{s}|_{2}\leq\frac{|\mathcal{L}_{s}|_{2}^{2}}{N}|\mathbf{B}|_{2}^{2}+\frac{1}{M}|\mathbf{C}|_{2}^{2},\quad|H_{h}|_{2}\leq\frac{1}{N}(C_{1}|\mathcal{L}_{s}|_{2}^{2}+2C_{2}|\mathcal{L}_{s}|_{2}+C_{3})|\mathbf{B}|_{2}^{2}.$
+where $|\mathcal{L}_{s}|_{2}$ is the $L^{2}$-operator norm of $\mathcal{L}_{s}$ and
+$C_{1}=\max_{\mathbf{z}}\phi^{2}(\mathbf{z}),\quad C_{2}=\max_{\mathbf{z}}|\phi(\mathbf{z})\phi_{t}(\mathbf{z})|,\quad\text{and}\quad C_{3}=\max_{\mathbf{z}}\phi_{t}^{2}(\mathbf{z}).$
+Remark 2.
+Notice the competition between the PDE residual points and the initial conditions points (reflected in $\mathbf{B}$ and $\mathbf{C}$) in the bound for $H_{s}$, where for $H_{h}$, it is all about the operator $\mathcal{L}_{s}$ and the scaling by $\phi$ and $\phi_{t}$.
+The bounds will become much more complex when $\mathcal{P}$ is not linear or $\phi$ also depends on $\mathbf{x}$.
+In order to give a better picture on the training decay rates, we provide a Neural Tangent Kernel (NTK) perspective on how hard constraints would work for reduce the spectral bias and training difficulties of PINNs on solving stiff time-dependent PDEs.
+To simplify the discussion, we will focus on the PDEs with periodic boundary.
+Let $\{\mathbf{z}_{i}=(t_{i},\mathbf{x}_{i})\}_{i=1}^{N}\subset(0,T]\times\Omega$ denote collocation (residual) points and $\{\bm{\zeta}_{j}=(0,\mathbf{x}_{j})\}_{j=1}^{M}$ denote initial-condition sample points.
+We define the two residual functions $r_{s}$ and $r_{h}$ as
+$r_{s}(\mathbf{z}_{i})=\partial_{t}u_{s}(\mathbf{z}_{i})+\mathcal{P}[u_{s}](\mathbf{z}_{i})-f(\mathbf{z}_{i}),$
 and
-$v(t,x)=\begin{bmatrix}t&1&\cos(\frac{2\pi}{P}x)&\sin(\frac{2\pi}{P}x)&\cos(\frac{2\pi}{P}2x)&\cdots&\cos(\frac{2\pi}{P}mx)&\sin(\frac{2\pi}{P}mx)\end{bmatrix}^{\top},$
-where $m>0$ is a positive integer (another hyper-parameter).
-With this composition, $u_{nn}$ will automatically become $P$-periodic dong2021method .
-For higher dimensional physical $\mathbf{x}$, similar transformation can be used.
-2.2 Mini-batching
+$r_{h}(\mathbf{z}_{i})=\phi(\mathbf{z}_{i})\partial_{t}u_{h}(\mathbf{z}_{i})+\mathcal{P}[\psi+\phi u_{h}](\mathbf{z}_{i})+\phi_{t}(\mathbf{z}_{i})u_{h}(\mathbf{z}_{i})-f(\mathbf{z}_{i})-\psi_{t}(\mathbf{z}_{i}),$
+where $u_{s}=u_{s}(t,\mathbf{x};\bm{\theta}_{s})$ (soft constrained) and $u_{h}=u_{h}(t,\mathbf{x};\bm{\theta}_{h})$ (hard constrained) are the two PINN approximations, which are automatically periodic due to the Fourier layer.
+The corresponding losses to train $u_{s}$ and $u_{h}$ are
+$\text{Loss}_{s}(u_{s}(\bm{\theta}_{s}))=\text{Loss}_{IC}(u_{s}(\bm{\theta}_{s}))+\text{Loss}_{PDE}(u_{s}(\bm{\theta}_{s}))$
+with
+$\begin{cases}\text{Loss}_{IC}(u_{s}(\bm{\theta}_{s}))&=\frac{1}{2M}\sum_{i=1}^{M}|u_{s}(\bm{\zeta}_{i};\bm{\theta}_{s})-u_{0}(\mathbf{x}^{IC}_{i})|^{2},\quad\bm{\zeta}_{i}=(0,\mathbf{x}^{IC}_{i}),\\ \text{Loss}_{PDE}(u_{s}(\bm{\theta}_{s}))&=\frac{1}{2N}\sum_{i=1}^{N}|r_{s}(\mathbf{z}_{i};\bm{\theta}_{s})|^{2},\end{cases}$
+and
+$\text{Loss}_{h}(u_{h}(\bm{\theta}_{h}))=\frac{1}{2N}\sum_{i=1}^{N}|r_{h}(\mathbf{z}_{i};\bm{\theta}_{h})|^{2}.$
+Both Losses can induce a continuous-time (in terms of $\tau$ not $t$ as $t$ represents the physical time) gradient flow systems as
+$\frac{\operatorname{d\!}\bm{\theta}_{s}(\tau)}{\operatorname{d\!}\tau}=-\nabla\text{Loss}_{s}(u_{s}(\bm{\theta}_{s}(\tau)))\quad\text{and}\quad\frac{\operatorname{d\!}\bm{\theta}_{h}(\tau)}{\operatorname{d\!}\tau}=-\nabla\text{Loss}_{h}(u_{h}(\bm{\theta}_{s}(\tau))),$
+with $\bm{\theta}_{s}(0)=\bm{\theta}_{h}(0)=\bm{\theta}_{0}$.
+Let the operators $\mathcal{L}_{s}[u]\coloneqq\partial_{t}u+\mathcal{P}[u]$ and $\mathcal{L}_{h}[u]\coloneqq\phi\partial_{t}u+\mathcal{P}[\psi+\phi u]+\phi_{t}u$.
+Then with NTK, we obtain
+$\begin{bmatrix}\frac{\operatorname{d\!}u_{s}(\bm{\zeta};\bm{\theta}_{s}(\tau))}{\operatorname{d\!}\tau}\\ \frac{\operatorname{d\!}\mathcal{L}_{s}[u_{s}](\mathbf{z};\bm{\theta}_{s}(\tau))}{\operatorname{d\!}\tau}\end{bmatrix}=-\begin{bmatrix}\mathbf{K}^{s}_{11}(\tau)&\mathbf{K}^{s}_{12}(\tau)\\ \mathbf{K}^{s}_{21}(\tau)&\mathbf{K}^{s}_{22}(\tau)\end{bmatrix}\cdot\begin{bmatrix}u_{s}(\bm{\zeta};\bm{\theta}_{s}(\tau))-u_{0}(\mathbf{x}^{IC})\\ \mathcal{L}_{s}[u_{s}](\mathbf{z};\bm{\theta}_{s}(\tau))-f(\mathbf{z})\end{bmatrix}$.
+with $\mathbf{K}^{s}_{11}(\tau)\in\mathbb{R}^{M\times M}$, $\mathbf{K}_{12}(\tau)\in\mathbb{R}^{M\times N}$ ($\mathbf{K}^{s}_{12}(\tau)=(\mathbf{K}^{s}_{21}(\tau))^{\top}$), and $\mathbf{K}^{s}_{22}(\tau)\in\mathbb{R}^{N\times N}$ and the entries are given as
+$\begin{cases}(\mathbf{K}^{s}_{11})_{i,j}(\tau)&=\langle\partial_{\bm{\theta}}u_{s}(\bm{\zeta}_{i};\bm{\theta}_{s}(\tau))\,,\partial_{\bm{\theta}}u_{s}(\bm{\zeta}_{j};\bm{\theta}_{s}(\tau))\rangle\\ (\mathbf{K}^{s}_{12})_{i,j}(\tau)&=\langle\partial_{\bm{\theta}}u_{s}(\bm{\zeta}_{i};\bm{\theta}_{s}(\tau))\,,\partial_{\bm{\theta}}\mathcal{L}_{s}[u_{s}](\mathbf{z}_{j};\bm{\theta}_{s}(\tau))\rangle\\ (\mathbf{K}^{s}_{22})_{i,j}(\tau)&=\langle\partial_{\bm{\theta}}\mathcal{L}_{s}[u_{s}](\mathbf{z}_{i};\bm{\theta}_{s}(\tau))\,,\partial_{\bm{\theta}}\mathcal{L}_{s}[u_{s}](\mathbf{z}_{j};\bm{\theta}_{s}(\tau))\rangle\\ \end{cases}.$
+Meanwhile, for $u_{h}$, we have
+$\frac{\operatorname{d\!}\mathcal{L}_{h}[u_{h}](\mathbf{z};\bm{\theta}_{h}(\tau))}{\operatorname{d\!}\tau}=-\mathbf{K}^{h}(\mathcal{L}_{h}[u_{h}](\mathbf{z};\bm{\theta}_{h}(\tau))-f(\mathbf{z})-\psi_{t}(\mathbf{z})),\quad\mathbf{K}^{h}\in\mathbb{R}^{M\times M},$
+with the entries of $\mathbf{K}^{h}$ being
+$(\mathbf{K}^{h})_{i,j}(\tau)=\langle\partial_{\bm{\theta}}\mathcal{L}_{h}[u_{h}](\mathbf{z}_{i};\bm{\theta}_{h}(\tau))\,,\partial_{\bm{\theta}}\mathcal{L}_{h}[u_{h}](\mathbf{z}_{j};\bm{\theta}_{h}(\tau))\rangle.$
+Setting $\mathbf{K}^{s}=\begin{bmatrix}\mathbf{K}^{s}_{11}&\mathbf{K}^{s}_{12}\\ \mathbf{K}^{s}_{21}&\mathbf{K}^{s}_{22}\end{bmatrix}$, and assuming $\mathbf{K}^{s}(\tau)\approx\mathbf{K}^{s}(0)=\mathbf{K}^{s}_{*}$ and $\mathbf{K}^{h}(\tau)\approx\mathbf{K}^{h}(0)=\mathbf{K}^{h}_{*}$, then the decays satisfy
+$\begin{bmatrix}\frac{\operatorname{d\!}u_{s}(\bm{\zeta};\bm{\theta}_{s}(\tau))}{\operatorname{d\!}\tau}\\ \frac{\operatorname{d\!}\mathcal{L}_{s}[u_{s}](\mathbf{z};\bm{\theta}_{s}(\tau))}{\operatorname{d\!}\tau}\end{bmatrix}-\begin{bmatrix}u_{0}(\mathbf{x}^{IC})\\ f(\mathbf{z})\end{bmatrix}=-e^{-\mathbf{K}^{s}_{*}\tau}\begin{bmatrix}u_{0}(\mathbf{x}^{IC})\\ f(\mathbf{z})\end{bmatrix}=-\mathbf{Q}_{s}^{\top}e^{-\bm{\Lambda}_{s}\tau}\mathbf{Q}_{s}\begin{bmatrix}u_{0}(\mathbf{x}^{IC})\\ f(\mathbf{z})\end{bmatrix}$,
+where $\mathbf{K}^{s}_{*}=\mathbf{Q}_{s}^{\top}\bm{\Lambda}_{s}\mathbf{Q}_{s}$; moreover
+$\frac{\operatorname{d\!}\mathcal{L}_{h}[u_{h}](\mathbf{z};\bm{\theta}_{h}(\tau))}{\operatorname{d\!}\tau}-(f(\mathbf{z})+\psi_{t}(\mathbf{z}))=-e^{\mathbf{K}^{h}_{*}\tau}(f(\mathbf{z})+\psi_{t}(\mathbf{z}))=-\mathbf{Q}_{h}^{\top}e^{-\bm{\Lambda}_{h}\tau}\mathbf{Q}(f(\mathbf{z})+\psi_{t}(\mathbf{z})),$
+where $\mathbf{K}^{h}_{*}=\mathbf{Q}_{h}^{\top}\bm{\Lambda}_{h}\mathbf{Q}_{h}$.
+One can see that the training decay of not using the hard-constrained transformation depends on both the eigenvalues of $\mathbf{K}^{s}_{11}$ and $\mathbf{K}^{s}_{22}$; whereas the training decay of the hard-constrained transformation depends only on $\mathbf{K}^{h}$, hence reducing the spectral bias.
+4 Examples
+In this section, we examine the hard-constrained PINN (HC-PINN) on a diverse set of prototypical stiff, time-dependent PDEs with periodic boundary conditions, aiming to demonstrate its robustness and flexibility in addressing a broad class of challenging problems.
+These equations often involve strong nonlinearities, stiffness, and parameter sensitivities, posing significant challenges for data-driven methods such as PINNs.
+Nonetheless, our results illustrate the capacity of the proposed framework to overcome these difficulties and accurately approximate the solutions across these varied regimes.
+To begin, we employ Latin hypercube sampling to generate collocation points across the full spatial-temporal domain, including its boundary.
+To facilitate training and better incorporate the initial condition into the solution structure, we adopt the following transformation:
+$\tilde{u}(t,\mathbf{x})=e^{-Ct}*u_{0}(\mathbf{x})+t*u_{nn}(t,\mathbf{x}),\quad\psi=e^{-Ct}*u_{0}(\mathbf{x}),\,\phi=t,\quad(t,\mathbf{x})\in[0,T]\times\bar{\Omega},$
+where $u_{0}(\mathbf{x})$ denotes the initial condition, and we test two values of $C\in\{0.1,1\}$, to assess how the decay rate affects learned dynamics.
+The transformed PINN $u_{nn}$ is trained to minimize a loss function that accounts for the PDE residuals, which now includes IC/BC.
+We present the results in four major categories, $(I)$ ablation study across $7$ different stiff PDEs where stiffness comes in different forms, such as sharp phase transition in Allen-Cahn, higher-order derivatives (forth order) in Cahn-Hilliard and Kuramoto-Sivashinsky, coupled outputs form Gray-Scott and Shrödinger systems, and high frequency modes in fast moving linear transport; $(II)$ comparison to other start-of-the-art PINNs such as the Causal PINN (Wang et al., 2022) and Enhanced RBA PINN (Anagnostopoulos et al., 2024) on Allen-Cahn, Cahn-Hilliard and Kuramoto-Sivashinsky; $(III)$ application of HC-PINNs on a $2D$ Allen-Cahn; $(IV)$ sensitivity analysis on the parameters $m$ for the Allen-Cahn equation with two different initial conditions.
+Causal training (MLP) Enhanced RBA HC-PINNs
+AC (Case I) $6.95\times 10^{-2}$ $2.62\times 10^{-3}$ $8.29\times 10^{-4}$
+AC (Case II) $1.78\times 10^{-2}$ $1.08\times 10^{-3}$ $2.14\times 10^{-4}$
+CH $3.49\times 10^{-1}$ $9.83\times 10^{-2}$ $5.61\times 10^{-4}$
+KS $2.72\times 10^{-1}$ $3.64\times 10^{-2}$ $5.37\times 10^{-4}$
+Table 1: Relative $L^{2}$ errors for the Allen–Cahn (AC), Cahn-Hilliard (CH) and Kuramoto–Sivashinsky (KS) equations obtained after $50k$ training iterations.
+All methods are implemented under the same $50k$ iteration budget for a fair comparison.
+Note that the reported errors for Causal training (Wang et al., 2022) and Enhanced RBA (Anagnostopoulos et al., 2024) differ from those in their original papers, where results were obtained with $300k$ iterations.
+Figure 1 provides a summary of our ablation study over $7$ different benchmarks; it shows the comparison between the reference exact solutions and our best HC-PINN predictions, together with the corresponding absolute errors; the benchmarks include Allen–Cahn, Cahn–Hilliard, Kuramoto–Sivashinsky, Gray–Scott, nonlinear Schrödinger, and linear advection equations (with different speeds $c=50,60$).
+Complementing these visual comparisons, Table 1 summarizes the comparison study between HC-PINN and two other PINN variants in terms of relative $L^{2}$ errors for Allen–Cahn, Cahn-Hilliard and Kuramoto–Sivashinsky equations under a strictly controlled setting: all methods are trained for $50k$ iterations, same data and parameter settings to ensure fairness.
+Notably, while prior works originally reported results with $300k$ iterations, HC-PINNs achieve substantially lower errors even with a much shorter training budget.
+Overall, these extensive experiments highlight the effectiveness and adaptability of our PINN scheme in solving a wide variety of stiff time-dependent PDEs.
+The results provide not only strong empirical validation but also a reliable and reproducible benchmark for future developments in this area.
+For full experimental details and visualizations, see Section B in the appendix.
+The results of solving the $2D$ Allen–Cahn equation in Figure 2 were obtained after training the network for $50k$ iterations using the initial condition $u(0,x,y)=\sin(4\pi x)\cos(4\pi y)$, with $\gamma_{1}=0.0001$ and $\gamma_{2}=1$ in Eq.
+5, and periodic boundary conditions.
+The HC-PINN accurately captures the solution evolution, yielding exceptionally small relative $L^{2}$ errors: $8.04\times 10^{-5}$ at $t=0.5$ and $1.33\times 10^{-4}$ at $t=1$.
+This demonstrates the high accuracy and stability of our method for 2D Allen–Cahn dynamics.
+At last, we perform a sensitivity study on the number of Fourier modes $m$ for the Allen–Cahn benchmarks to assess how this key hyperparameter influences the accuracy of HC-PINNs.
+The results demonstrate that accuracy is highly dependent on spectral resolution: while small $m$ yields stable and low errors, performance deteriorates significantly once $m$ exceeds a critical threshold.
+Exact $u(x,y)$ Predicted $u(x,y)$ Error
+Figure 2: Reference, HC-PINN predicted solution, and absolute error of the 2D Allen Cahn equation at different time snapshots (a) $t=0$ , (b) $t=0.5$ , (c) $t=1$.
+Figure 3: Sensitivity of HC-PINNs to the number of Fourier modes $m$ in the Allen–Cahn equations. The plot shows the relative $L^{2}$ error as a function of $m$.
+5 Conclusion
+We provide a comprehensive study of hard-constraints and self-adaptive loss weights to study the stabilization of PINN training on solving stiff time-dependent PDEs.
+Our comprehensive study includes an ablation study of seven distinct types of PDES, comparison study, a theoretical analysis, an application to $2D$ Allen Chan, and a sensitive analysis of $m$ (the number of Fourier modes).
+From the ablation study, we found out that the transformation provided by hard constraint significantly alleviates the training challenges typically associated with stiff time-dependent PDEs.
+Our ablation study demonstrates effectiveness across seven distinct types of such PDEs, encompassing scalar-valued and vector-valued (including complex-valued) outputs.
+This approach not only enhances training efficiency but also improves the accuracy and stability of the learned solutions, particularly in the presence of stiffness, where conventional PINNs often struggle.
+We provide theoretical insights using NTK into how hard-constrained transformation is reducing the training difficulties, especially in terms of reducing the spectral bias.
+The comparison study focuses on comparing our HC-PINN to two other state-of-the-art PINN variants on established examples.
+Future work includes extending the hard-cosntraint framework to more two- and three-dimensional PDEs, which pose additional challenges in both architecture design and computational cost.
+Ongoing research also focuses on identifying optimal transformations $\psi$ and $\phi$, which govern the embedding of IC and BC into the network structure, with the goal of further enhancing performance and generalizability.
+Appendix A Methodology: Additional Details
+In this section, we discuss the additional methods used in the ablation study, including mini-batching and self-adaptive loss weights.
+A.1 Mini-batching
 We have eliminated the IC/BC losses, but the transformed PDE loss might create additional difficulties in training due to the introduction of IC/BC functions into the PDE.
@@
 Sampling mini-batches from collocation points also helps stabilize and accelerate optimization.
-So we consider mini-batching in the training for decreasing the PDE residual loss.
-As pointed out in wight2020solving , the mini-batching technique is similar to time-marching resampling.
-But with mini-batching, one does not employ an explicit time grid.
-2.3 Self-Adaptive Physics-Informed Neural Networks
-The proper choice of penalty term $\lambda$ can be updated from derived formula as mentioned in wang2020and .
+So we consider mini-batching in the training for the PDE residual loss.
+As pointed out in (Wight & Zhao, 2020), the mini-batching technique is similar to time-marching resampling but without an explicit time grid.
+A.2 Self-Adaptive PINNs
+The proper choice of penalty term $\lambda$ can be updated from derived formula as mentioned in (Wang et al., 2020b).
 However, data-driven discovery of $\lambda$ is possible.
-Outline in mcclenny2020self , the loss is further modified as
-$\text{Data Loss}(\lambda_{IC,i}\}_{i=1}^{N_{IC}},\{\lambda_{BC,i}\}_{i=1}^{N_{BC}})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}m(\lambda_{IC,i})|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
-$\quad+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}m(\lambda_{BC,i})|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2}$
-Physical $\text{Physical Loss}(\{\lambda_{PDE,i}\}_{i=1}^{N_{PDE}})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}m(\lambda_{PDE,i})|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}$
-The set of loss weights, $\{\lambda_{IC,i}\}_{i=1}^{N_{IC}}\cup\{\lambda_{BC,i}\}_{i=1}^{N_{BC}}\cup\{\lambda_{PDE,i}\}_{i=1}^{N_{PDE}}$, are designed in a way to increase when needed during any training epoch in order to combat the spectral bias.
+Outline in (McClenny & Braga-Neto, 2020), the loss is further modified as
+$\text{Data Loss}(\{\lambda^{IC}_{i}\}_{i=1}^{N_{IC}},\{\lambda^{BC}_{i}\}_{i=1}^{N_{BC}})$
+$=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}F_{mask}(\lambda^{IC}_{i})|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
+$+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}F_{mask}(\lambda^{BC}_{j})|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2},$
+and
+$\text{Physical Loss}(\{\lambda^{CL}_{i}\}_{i=1}^{N_{CL}})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}F_{mask}(\lambda^{CL}_{i})|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}.$
+The set of loss weights, $\{\lambda^{IC}_{i}\}_{i=1}^{N_{IC}}\cup\{\lambda^{BC}_{i}\}_{i=1}^{N_{BC}}\cup\{\lambda^{CL}_{i}\}_{i=1}^{N_{CL}}$, are designed in a way to increase when needed during any training epoch in order to combat the spectral bias.
 Adaptive methods are essential to ensure that the neural network accurately addresses the challenging spots in the solutions of “stiff” PDEs.
-In our experiments, we employ self-adaptive PINNs as proposed by mcclenny2020self to train PINNs adaptively.
+In our experiments, we employ self-adaptive PINNs as proposed by (McClenny & Braga-Neto, 2020) to train PINNs adaptively.
 This approach involves fully trainable adaptation weights applied individually to each training point, allowing the neural network to independently identify and focus on the challenging regions of the solution.
-2.4 Performance Measures
+A.3 Performance Measures
 Let $\{x_{i},t_{i}\}_{i=1}^{N}$ be a collection of $N$ data points where we evaluate both the reference solution $u(x,t)$ and the neural network output $\mathcal{U}(x,t)$.
 To measure the accuracy of the trained model, we compute the relative $L_{1}$ norm $\mathcal{E}_{1}$ and relative $L_{2}$ norm $\mathcal{E}_{2}$, defined by:
-$\mathcal{E}_{1}=\frac{\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|}{\sum_{i=1}^{N}|u(x_{i},t_{i})|},\quad\quad\quad\mathcal{E}_{2}=\frac{\Big{(}\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|^{2}\Big{)}^{1/2}}{\sqrt{\sum_{i=1}^{N}|u(x_{i},t_{i})|^{2}}}.$
-These metrics represent the normalized differences between the model predictions and the ground truth values, providing a quantitative evaluation of model performance.
-3 Numerical Experiments
-In this section, we examine the proposed PINN scheme on a diverse set of prototypical stiff time-dependent PDEs with periodic boundary conditions, aiming to demonstrate its robustness and flexibility in handling a broad class of challenging equations.
-These benchmarks exhibit a wide range of physical and mathematical characteristics: from stiff dynamics (e.g., Kuramoto–Sivashinsky), to multi-species reaction–diffusion systems (e.g., Gray–Scott and Belousov–Zhabotinsky), and nonlinear dispersive waves (e.g., nonlinear Schrödinger equation).
-Such equations often involve strong nonlinearities, stiffness, and parameter sensitivities, posing significant challenges for data-driven methods such as PINNs.
-Nonetheless, our results illustrate the capacity of the proposed framework to overcome these difficulties and accurately approximate the solutions across these varied regimes.
-Detailed experimental results are presented in Section 3.1.
-To begin, we employ Latin Hypercube Sampling to generate collocation points across the full spatio-temporal domain, including its boundary.
-To facilitate training and better incorporate the initial condition into the solution structure, we adopt the following transformation:
-$\tilde{u}(t,\mathbf{x})=\exp(-Ct)*u_{0}(\mathbf{x})+t*u_{nn}(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\bar{\Omega};\quad\psi=\exp(-Ct)*u_{0}(\mathbf{x}),\,\phi=t.$
-where $u_{0}(\mathbf{x})$ denotes the initial condition, and we test two values of $C\in\{0.1,1\}$, to assess how the decay rate affects learned dynamics.
-The transformed PINN $u_{nn}$ is trained to minimize a loss function that accounts for the PDE residuals, which now includes IC/BC.
-To rigorously assess the accuracy of our results, we use Chebfun, a spectral-style computing system in MATLAB that enables highly accurate function approximations.
-This allows us to make precise comparisons between the ground-truth solutions and the PINN-generated solutions.
-For each benchmark, we not only present the main numerical results but also carry out detailed ablation studies to evaluate the contribution of each proposed technique.
-In each ablation experiment, we deactivate one component of the methodology at a time, keeping the remaining elements active under the same hyper-parameter configurations, to measure the relative $L^{2}$ error and relative $L^{1}$ error.
-The neural network architecture used throughout consists of 7 hidden layers with 32 neurons per layer and tanh activation functions.
-The initial weights are fixed across all ablation runs for consistency.
-Model training proceeds in two stages: first using the Adam optimizer kingma2014adam with a learning rate of $10^{-3}$ for $50,000$ steps, followed by fine-tuning via the L-BFGS-B optimizer with a learning rate of $1$ for $5,000$ steps.
-Each experiment uses $16,384$ interior collocation points, along with $128$ initial and $128$ boundary data points sampled via LHS.
-For PDE systems (e.g., Gray–Scott), we further apply a mini-batching strategy with $5$ batches and a batch size of $3,280$ to enhance training stability.
-Overall, these extensive experiments highlight the effectiveness and adaptability of our PINN scheme in solving a wide variety of stiff time-dependent PDEs.
-The results provide not only strong empirical validation but also a reliable and reproducible benchmark for future developments in this area.
-For full experimental details and visualizations, see Section 3.1.
-3.1 Allen-Cahn Equation $I$
-The Allen-Cahn type PDE is a classical phase-field model that has been widely used to study phase separation phenomena bazant2017thermodynamic .
-It has numerous practical applications across a range of fields, including materials science allen1979microscopic ; shen2010numerical , biological systems hyman2014liquid ; takatori2015towards , and electrochemical systems horstmann2013precipitation ; tian2015electrochemically .
-We begin with the one-dimensional Allen-Cahn equation with periodic boundary conditions, a representative example that presents challenges for traditional PINN models, formulated as follows:
-$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}u^{3}-\gamma_{2}u=0,\quad(t,x)\in(0,T)\times(a,b),$ (3)
-$u(0,x)=u_{0}(x),\quad x\in[a,b],$
-$u(t,a)=u(t,b),\quad t\in[0,T],$
-$u_{x}(t,a)=u_{x}(t,b),\quad t\in[0,T],$
+$\small\mathcal{E}_{1}=\frac{\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|}{\sum_{i=1}^{N}|u(x_{i},t_{i})|},\quad\mathcal{E}_{2}=\frac{\sqrt{\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|^{2}}}{\sqrt{\sum_{i=1}^{N}|u(x_{i},t_{i})|^{2}}}.$
+These metrics measure the discrepancy between the predicted and reference solutions in average (energy-based) and absolute (mass-based) senses, respectively.
+The relative form normalizes the error with respect to the reference solution, enabling fair comparisons across different scales and variables.
+To rigorously assess the accuracy of our results, we implemented a semi-implicit spectral method in Matlab.
+This allows us to make precise comparisons between the ground-truth solutions $u(x_{i},t_{i})$ and the PINN-generated solutions $\mathcal{U}(x_{i},t_{i})$.
+A.4 Proofs of the Theorems
+We present the proofs for the two aforementioned theorem.
+Theorem (Well Conditioning).
+The transformation is well defined when $\phi$ and $\psi$ is $C^{1}$ in time and have $K^{th}$ order partial derivatives w.r.t $\mathbf{x}$.
+Moreover, $\phi_{t}(0,\mathbf{x})\neq 0$.
+Then $u_{nn}$ will satisfy a new PDE
+$\begin{cases}&\phi\partial_{t}u_{nn}+\mathcal{P}[\psi+\phi u_{nn}]+\phi_{t}u_{nn}=f-\psi_{t},\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ &u_{nn}(0,\mathbf{x})=\frac{u_{t}(0,\mathbf{x})-\psi_{t}(0,\mathbf{x})}{\phi_{t}(0,\mathbf{x})},\quad\mathbf{x}\in\bar{\Omega}.\end{cases}$
+Proof.
+Since $\mathcal{L}_{s}[u]=\partial_{t}u+\mathcal{P}[u]$ is applied to the transformation $u=\psi+\phi u_{nn}$ and $u_{nn}$ is constructed with hyper tangent as the activation function, we only need $\psi$ and $\phi$ to satisfy the required regularity from $\mathcal{L}_{s}$ so that the operator can still be applied, hence $\psi$ and $\phi$ have to have first-order time derivative and $K^{th}$-order partial derivatives w.r.t to $\mathbf{x}$, we further require them to be continuous in order to allow classical solutions to exist for $\mathcal{L}_{s}$.
+Meanwhile from the transformation we have
+$u_{nn}(t,\mathbf{x})=\frac{u(t,\mathbf{x})-\psi(t,\mathbf{x})}{\phi(t,\mathbf{x})},\quad\text{for $(t,\mathbf{x})\in(0,T]\times\Omega$.}$
+As $t\rightarrow 0$, we have $\psi(0,\mathbf{x})=u_{0}(\mathbf{x})$ and $\phi(0,\mathbf{x})=0$, hence by L’Hôpital’s rule
+$\lim_{t\rightarrow 0}u_{nn}(t,\mathbf{x})=\lim_{t\rightarrow 0}\frac{u(t,\mathbf{x})-\psi(t,\mathbf{x})}{\phi(t,\mathbf{x})}=\lim_{t\rightarrow 0}\frac{u_{t}(t,\mathbf{x})-\psi_{t}(t,\mathbf{x})}{\phi_{t}(t,\mathbf{x})},$
+when $\phi_{t}(0,\mathbf{x})\neq 0$, then limit is defined and we can have
+$u_{nn}(0,\mathbf{x})=\lim_{t\rightarrow 0}u_{nn}(t,\mathbf{x})=\frac{u_{t}(0,\mathbf{x})-\psi_{t}(0,\mathbf{x})}{\phi_{t}(0,\mathbf{x})}.$
+Moreover
+$\mathcal{L}_{s}(\psi+\phi u_{nn})=\partial_{t}(\psi+\phi u_{nn})+\mathcal{P}[\psi+\phi u_{nn}]=\phi\partial_{t}u_{nn}+\mathcal{P}[\psi+\phi u_{nn}]+\phi_{t}u_{nn}+\psi_{t},$
+hence the new PDE for $u_{nn}$. ∎
+Next, the theorem on Hessian bounds.
+Theorem.
+Assume that $\mathcal{P}$ is linear and $\phi(t,\mathbf{x})=\phi(t)\neq 0$ for $t>0$, then for a one-hidden layer network, we have for a network $u(\mathbf{y};\theta)$ with $\bm{\theta}\in\mathbb{R}^{D}$ being the network parameters, the following relationship is satisfied
+$u(\mathbf{y};\bm{\theta})=\mathbf{h}^{\top}(\mathbf{y})\bm{\theta},\quad\mathbf{h}=\begin{bmatrix}h_{1}(\mathbf{y})&\cdots&\mathbf{h}_{D}(\mathbf{y})\end{bmatrix}^{\top}$,
+where $\mathbf{y}=\mathbf{z}=(t,\mathbf{x})\in(0,T]\times\Omega$ or $\mathbf{y}=\bm{\zeta}=(0,\mathbf{x}^{IC})\in\{0\}\times\bar{\Omega}$. Therefore the two losses become
+$\text{Loss}_{s}(u(\mathbf{y};\bm{\theta}))$ $=\frac{1}{2N}|\mathbf{A}\bm{\theta}-\mathbf{f}|_{\ell^{2}(\mathbb{R}^{N})}^{2}+\frac{1}{2M}|\mathbf{C}\bm{\theta}-\mathbf{u}_{0}|_{\ell^{2}(\mathbb{R}^{M})}^{2},\quad\text{(soft constrained)}$
+$\text{Loss}_{h}(u(\mathbf{z};\bm{\theta}))$ $=\frac{1}{2N}|\bm{\Lambda}_{\phi}\mathbf{A}\bm{\theta}+\bm{\Lambda}_{\phi_{t}}\mathbf{B}\bm{\theta}-\tilde{\mathbf{f}}|_{\ell^{2}(\mathbb{R}^{N})}^{2},\quad\text{(hard constrained)}$
+where
+$\mathbf{A}=\begin{bmatrix}\mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{1})]\\ \vdots\\ \mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{N})]\end{bmatrix},\quad\mathbf{B}=\begin{bmatrix}\mathbf{h}^{\top}(\mathbf{z}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\mathbf{z}_{N})\end{bmatrix},\quad\text{and}\quad\mathbf{C}=\begin{bmatrix}\mathbf{h}^{\top}(\bm{\zeta}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\bm{\zeta}_{M})\end{bmatrix}$,
+where $\mathbf{x}^{IC}_{i}\in\bar{\Omega}$; and for the diagonal matrices $\bm{\Lambda}_{\phi}$ and $\bm{\Lambda}_{\phi_{t}}$
+$\bm{\Lambda}_{\phi}=\text{diag}(\phi(\mathbf{z}_{1}),\cdots,\phi(\mathbf{z}_{N}))\quad\text{and}\quad\bm{\Lambda}_{\phi_{t}}=\text{diag}(\phi_{t}(\mathbf{z}_{1}),\cdots,\phi_{t}(\mathbf{z}_{N})),$
+and the vectors
+$\mathbf{f}=\begin{bmatrix}f(\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})\end{bmatrix},\quad\tilde{\mathbf{f}}=\begin{bmatrix}f(\mathbf{z}_{1})-\mathcal{L}_{s}[\psi](\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})-\mathcal{L}_{s}[\psi](\mathbf{z}_{N})\end{bmatrix},\quad\text{and}\quad\mathbf{u}_{0}=\begin{bmatrix}u_{0}(\mathbf{x}_{1}^{IC})\\ \vdots\\ u_{0}(\mathbf{x}_{M}^{IC})\end{bmatrix}$.
+Therefore the Hessians of the losses are
+$H_{s}(\bm{\theta})$ $=\frac{1}{N}\mathbf{A}^{\top}\mathbf{A}+\frac{1}{M}\mathbf{C}^{\top}\mathbf{C},$
+$H_{h}(\bm{\theta})$ $=\frac{1}{N}(\mathbf{A}^{\top}\bm{\Lambda}_{\phi}^{2}\mathbf{A}+2\mathbf{A}^{\top}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\phi_{t}}\mathbf{B}+\mathbf{B}^{\top}\bm{\Lambda}_{\phi_{t}}^{2}\mathbf{B}),$
+Then the Hessians are bounded as
+$|H_{s}|_{2}$ $\leq\frac{|\mathcal{L}_{s}|_{2}^{2}}{N}|\mathbf{B}|_{2}^{2}+\frac{1}{M}|\mathbf{C}|_{2}^{2},$
+$|H_{h}|_{2}$ $\leq\frac{1}{N}(C_{1}|\mathcal{L}_{s}|_{2}^{2}+2C_{2}|\mathcal{L}_{s}|_{2}+C_{3})|\mathbf{B}|_{2}^{2}.$
+where $|\mathcal{L}_{s}|_{2}$ is the $L^{2}$-operator norm of $\mathcal{L}_{s}$ and
+$C_{1}=\max_{\mathbf{z}}\phi^{2}(\mathbf{z}),\quad C_{2}=\max_{\mathbf{z}}|\phi(\mathbf{z})\phi_{t}(\mathbf{z})|,\quad\text{and}\quad C_{3}=\max_{\mathbf{z}}\phi_{t}^{2}(\mathbf{z}).$
+Proof.
+Once we have $u(\mathbf{y};\bm{\theta})=\mathbf{h}^{\top}(\mathbf{y})\bm{\theta}$, then the loss from the soft-constrained PINN can be re-written as
+$\text{Loss}_{s}(u(\bm{\theta}))=\frac{1}{2N}\sum_{i=1}^{N}|(\mathcal{L}_{s}[u(\cdot;\bm{\theta})]-f)(\mathbf{z}_{i})|^{2}+\frac{1}{2M}\sum_{j=1}^{M}|u(\bm{\zeta}_{i};\bm{\theta})-u_{0}(\mathbf{x}_{i}^{IC})|^{2},$
+If we define the matrices
+$\mathbf{A}=\begin{bmatrix}\mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{1})]\\ \vdots\\ \mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{N})]\end{bmatrix}\quad\text{and}\quad\mathbf{C}=\begin{bmatrix}\mathbf{h}^{\top}(\bm{\zeta}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\bm{\zeta}_{M})\end{bmatrix}$,
+and the vectors
+$\mathbf{f}=\begin{bmatrix}f(\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})\end{bmatrix}\quad\text{and}\quad\mathbf{u}_{0}=\begin{bmatrix}u_{0}(\mathbf{x}_{1}^{IC})\\ \vdots\\ u_{0}(\mathbf{x}_{M}^{IC})\end{bmatrix}$.
+then $\text{Loss}_{s}(u(\bm{\theta}))=\frac{1}{2N}|\mathbf{A}\bm{\theta}-\mathbf{f}|_{\ell^{2}(\mathbb{R}^{N})}^{2}+\frac{1}{2M}|\mathbf{C}\bm{\theta}-\mathbf{u}_{0}|_{\ell^{2}(\mathbb{R}^{M})}^{2}$.
+Similarly, for the loss
+$\text{Loss}_{h}(u(\bm{\theta}))=\frac{1}{2N}\sum_{i=1}^{N}|(\mathcal{L}_{s}[\psi(\cdot)+\phi(\cdot)u(\cdot;\bm{\theta})]-f)(\mathbf{z}_{i})|^{2},$
+Since
+$\mathcal{L}_{s}[\psi+\phi u]=\phi(u_{t}+\mathcal{P}[u])+\phi_{t}u+\psi_{t},\quad\text{$\mathcal{P}$ is linear and $\phi(t,\mathbf{x})=\phi(t)$}.$
+Hence with
+$\mathbf{B}=\begin{bmatrix}\mathbf{h}^{\top}(\mathbf{z}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\mathbf{z}_{N})\end{bmatrix}\quad\text{and}\quad\tilde{\mathbf{f}}=\begin{bmatrix}f(\mathbf{z}_{1})-\mathcal{L}_{s}[\psi](\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})-\mathcal{L}_{s}[\psi](\mathbf{z}_{N})\end{bmatrix}$,
+and the two diagonal matrices
+$\bm{\Lambda}_{\phi}=\text{diag}(\phi(\mathbf{z}_{1}),\cdots,\phi(\mathbf{z}_{N}))\quad\text{and}\quad\bm{\Lambda}_{\phi_{t}}=\text{diag}(\phi_{t}(\mathbf{z}_{1}),\cdots,\phi_{t}(\mathbf{z}_{N})),$
+we have $\text{Loss}_{h}(u(\bm{\theta}))=\frac{1}{2N}|\bm{\Lambda}_{\phi}\mathbf{A}\bm{\theta}+\bm{\Lambda}_{\phi_{t}}\mathbf{B}\bm{\theta}-\tilde{\mathbf{f}}|_{\ell^{2}(\mathbb{R}^{N})}^{2}$.
+Then the two gradients are
+$\nabla_{\bm{\theta}}\text{Loss}_{s}(u(\bm{\theta}))$ $=\frac{1}{N}(\mathbf{A}^{\top}\mathbf{A}\bm{\theta}-\mathbf{A}^{\top}\mathbf{f})+\frac{1}{M}(\mathbf{C}^{\top}\mathbf{C}\bm{\theta}-\mathbf{C}^{\top}\mathbf{u}_{0}),$
+[…]
+Differentiating w.r.t to $\bm{\theta}$ again, we obtain the Hessians as
+$H_{s}(\bm{\theta})$ $=\frac{1}{N}\mathbf{A}^{\top}\mathbf{A}+\frac{1}{M}\mathbf{C}^{\top}\mathbf{C},$
+$H_{h}(\bm{\theta})$ $=\frac{1}{N}(\mathbf{A}^{\top}\bm{\Lambda}_{\phi}^{2}\mathbf{A}+2\mathbf{A}^{\top}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\phi_{t}}\mathbf{B}+\mathbf{B}^{\top}\bm{\Lambda}_{\phi_{t}}^{2}\mathbf{B}),$
+Recall that $\mathcal{L}_{s}[u]=u_{t}+\mathcal{P}[u]$, hence by the property of an $L^{2}$ function/operator norm on $\mathcal{L}_{s}$, we have
+$|\mathcal{L}_{s}[u]|_{2}\leq|\mathcal{L}_{s}|_{2}|u|_{2},\quad\text{This $|\cdot|_{2}$ is a norm on functions.}$
+Then we have $|\mathbf{A}^{\top}\mathbf{A}|_{2}\leq|\mathcal{L}_{s}|_{2}^{2}|\mathbf{B}|_{2}^{2}$ (Here $|\mathbf{A}|_{2}$ is a matrix norm). And if we define
+$C_{1}=\max_{\mathbf{z}}\phi^{2}(\mathbf{z}),\quad C_{2}=\max_{\mathbf{z}}|\phi(\mathbf{z})\phi_{t}(\mathbf{z})|,\quad\text{and}\quad C_{3}=\max_{\mathbf{z}}\phi_{t}^{2}(\mathbf{z}).$
+We can have
+$|\mathbf{A}^{\top}\bm{\Lambda}_{\phi}^{2}\mathbf{A}|_{2}$ $\leq C_{1}|\mathbf{A}^{\top}\mathbf{A}|_{2}\leq|\mathcal{L}_{s}|_{2}^{2}|\mathbf{B}|_{2}^{2},$
+$|\mathbf{A}^{\top}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\phi_{t}}\mathbf{A}|_{2}$ $\leq C_{2}|\mathbf{A}^{\top}\mathbf{B}|\leq C_{2}|\mathcal{L}_{s}|_{2}|\mathbf{B}|_{2}^{2},$
+$|\mathbf{B}^{\top}\bm{\Lambda}_{\phi_{t}}\mathbf{B}|_{2}$ $\leq C_{3}|\mathbf{B}^{\top}\mathbf{B}|_{2}\leq C_{3}|\mathbf{B}|_{2}^{2}$
+Putting them back in, we obtain the final bounds
+$|H_{s}|_{2}$ $\leq\frac{|\mathcal{L}_{s}|_{2}^{2}}{N}|\mathbf{B}|_{2}^{2}+\frac{1}{M}|\mathbf{C}|_{2}^{2},$
+$|H_{h}|_{2}$ $\leq\frac{1}{N}(C_{1}|\mathcal{L}_{s}|_{2}^{2}+2C_{2}|\mathcal{L}_{s}|_{2}+C_{3})|\mathbf{B}|_{2}^{2}.$
+Appendix B Examples: Additional Details
+We provide additional details for the ablation study, comparison study, $2D$ application and sensitivity study of $m$ here.
+We also explain why each PDE or PDE system is chosen due to their particular difficulties for solvers in each sub-section.
+B.1 Implementation Details
+All experiments were implemented in PyTorch ($v2.2$) and executed on both CUDA-enabled NVIDIA GPUs and Apple Silicon devices via the MPS backend.
+Our code runs efficiently across Linux (Ubuntu $22.04$), macOS (Sonoma), and Windows $11$ platforms.
+Primary training was conducted on an NVIDIA $A100$ GPU ($40$ GB memory) and a MacBook Pro with Apple M1 Pro chip ($16$-core GPU, $32$ GB unified memory).
+For each benchmark, we used a fully connected neural network with $7$ hidden layers, each containing $32$ neurons and tanh activation functions, to represent the latent solution functions.
+The same initial weights were used across all ablation experiments for consistency.
+Model training was carried out using a two-stage strategy: first, we trained using the Adam optimizer with an initial learning rate of $10^{-3}$ for $50,000$ steps.
+This was followed by fine-tuning with the L-BFGS-B optimizer using a learning rate of $1$ for $5,000$ steps.
+We did not perform hyperparameter sweeps, and the total number of training epochs was fixed across all experiments.
+This paper does not rely on any datasets.
+We used Latin Hypercube Sampling (LHS) to generate a fixed set of $16,384$ interior collocation points, along with $128$ initial data points and $128$ boundary data points.
+This same sampling configuration was used consistently across all experiments to ensure comparability.
+For PDE systems, we employed a mini-batching strategy with $5$ batches of size $3,280$ to improve memory efficiency during training.
+Ablation studies were conducted by disabling one component of our method at a time, while keeping all other settings fixed, to evaluate its contribution using relative $L^{2}$ and relative $L^{1}$ error metrics, which measure the discrepancy between the predicted and reference solutions in average (energy-based) and absolute (mass-based) senses, respectively.
+The relative form normalizes the error with respect to the reference solution, enabling fair comparisons across different scales and variables.
+We also compare HC-PINN with causal training (MLP) and enhanced RBA using the same data and a total of 50k iterations, instead of the 300k iterations used in the original works.
+For the sensitivity study, experiments with different $m$ values are conducted on the same points, using the same number of iterations and the same architecture.
+For 2D training, we use 4,0964{,}096 initial collocation points, 10,00010{,}000 residual collocation points per segment, and $800$ boundary points per segment.
+The architecture is the same as in the 1D training.
+B.2 Allen-Cahn PDE
+The Allen-Cahn type PDE is a classical phase-field model that has been widely used to study phase separation phenomena Bazant (2017).
+It has numerous practical applications across a range of fields, including materials science Allen & Cahn (1979); Shen & Yang (2010), biological systems Hyman et al.
+(2014); Takatori & Brady (2015), and electrochemical systems Horstmann et al.
+(2013); Tian et al.
+(2015).
+We begin with the $1$-dim Allen-Cahn equation with periodic boundary conditions formulated as follows:
+$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}(u^{3}-u)=0,\,(t,x)\in(0,T]\times(a,b),$ (5)
+$u(0,x)=u_{0}(x),\,x\in[a,b],\quad u(t,a)=u(t,b),\,t\in[0,T],$
 where $\gamma_{1},\gamma_{2}>0,T>0,a<b$ are prescribed constants.
 As $\gamma_{2}$ increases, the transition interface of the solutions is sharper, which makes it harder to solve the AC equation numerically.
-We first demonstrate the effectiveness of our scheme by testing on Allen-Cahn equation with the following initial condition and parameters, $u_{0}(x)=x^{2}\cos(\pi x)$, $T=1$, $a=-1$, $b=1$, $\gamma_{1}=0.001$, and $\gamma_{2}=5$.
-Figure 1: Exact solution of the 1D Allen-Cahn Equation (Case I) with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $1.18\times 10^{-3}$ and the relative $L^{1}$ error is $5.34\times 10^{-4}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 2: Best solutions of the Allen-Cahn Equation (Case I) $\gamma_{2}=5$.
-The resulting relative $L^{2}$ error is $1.18\times 10^{-3}$ and the relative $L^{1}$ error is $5.34\times 10^{-4}$.
-The hyper-parameter configuration is specified in Section 3.
-For this example, we first test the baseline PINN approach raissi2019physics .
+B.2.1 Case II
+We first tested on Allen-Cahn equation with the following, $u_{0}(x)=x^{2}\cos(\pi x)$, $T=1$, a=1a=-1, $b=1$, $\gamma_{1}=0.001$, and $\gamma_{2}=5$.
+For this example, we first test the baseline PINN approach Raissi et al.
+(2019).
 As shown in Table 2, the baseline PINN alone fails to accurately solve the Allen-Cahn equation, motivating the development of a more effective training pipeline.
@@
 The results are summarized in Table 2.
-The full scheme, including self-adaptive loss balancing, achieves the best performance, reaching a relative $L^{2}$ error of $1.18\times 10^{-3}$ and relative $L^{1}$ error of $5.34\times 10^{-4}$.
+The full scheme, including self-adaptive loss balancing, achieves the best performance, reaching a relative $L^{2}$ error of $8.29\times 10^{-4}$ and relative $L^{1}$ error of $5.34\times 10^{-4}$.
 In particular, using the faster-decaying initial condition ($IC1$) leads to better results than the slower-decaying case ($IC2$), suggesting that the initial condition plays a more critical role during the early stages of training but becomes less dominant over time.
@@
 Removing any single part generally leads to performance degradation, with the most severe impact observed when the initial condition is omitted, resulting in a relative error close to $0.99$ in both $L^{1}$ and $L^{2}$ norms.
-Figure 1 illustrates the exact solution, the network prediction, and the absolute error distribution.
-As seen, the predicted solution matches the ground truth remarkably well, with errors concentrated at negligible levels.
 Figure 4 presents cross-sectional comparisons at different times, further confirming the excellent agreement between the predicted and true solutions across the entire time horizon.
 These results provide strong evidence for the effectiveness of the proposed training strategy and underscore the critical role of each pipeline component.
-Table 1: Allen-Cahn Equation (Casse $I$ and Case $II$): Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
-Ablation Settings Performance - Case $I$ Performance - Case $II$
+Ablation Settings Case II Case IIII
 PBC IC1 IC2 SA Rel. $L^{2}$ error Rel. $L^{1}$ error Rel. $L^{2}$ error Rel. $L^{1}$ error
-$1.18\times 10^{-3}$ $5.34\times 10^{-4}$ $3.53\times 10^{-4}$ $2.14\times 10^{-4}$
+$8.29\times 10^{-4}$ $5.34\times 10^{-4}$ $3.53\times 10^{-4}$ $2.14\times 10^{-4}$
 $7.64\times 10^{-2}$ $1.66\times 10^{-2}$ $7.04\times 10^{-4}$ $3.26\times 10^{-4}$
@@
 $2.86\times 10^{-2}$ $8.20\times 10^{-3}$ $9.80\times 10^{-4}$ $3.91\times 10^{-4}$
-$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
-$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
+$99\times 10^{-1} 9.9$$9\times 10^{-1} 9.9$$9\times 10^{-1} 9.9$$99\times 10^{-1}$
+$99\times 10^{-1} 9.9$$9\times 10^{-1} 9.9$$9\times 10^{-1} 9.9$$99\times 10^{-1}$
 $1.05\times 10^{-2}$ $2.65\times 10^{-3}$ $5.11\times 10^{-2}$ $1.21\times 10^{-2}$
@@
 $5.12\times 10^{-1}$ $3.18\times 10^{-1}$ $4.02\times 10^{-1}$ $1.42\times 10^{-1}$
-$9.99\times 10^{-1}$ $9.98\times 10^{-1}$ $5.11\times 10^{-1}$ $3.28\times 10^{-1}$
-3.2 Allen-Chan, Case $II$
-For our second example, we investigate the performance of PINNs in scenarios involving sharper moving interfaces.
-Specifically, we consider a modified Allen–Cahn equation with more oscillatory initial data:
-$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}u^{3}-\gamma_{2}u=0,\quad(t,x)\in(0,1)\times(-1,1),$ (4)
-$u(0,x)=x^{2}\sin(2\pi x),\quad x\in[-1,1],$
-$u(t,-1)=u(t,1),\quad t\in[0,1],$
-$u_{x}(t,-1)=u_{x}(t,1),\quad t\in[0,1],$
-where $\gamma_{1}=0.001,\gamma_{2}=4$.
+$99\times 10^{-1} 9.9$$98\times 10^{-1}$ $5.11\times 10^{-1}$ $3.28\times 10^{-1}$
+Table 2: Allen-Cahn (Casse II and Case IIII): Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
+B.2.2 Case IIII
+For the second case, we change the IC to have highly oscillatory data by $u(0,x)=x^{2}\sin(2\pi x)$. Other parameters are $T=1$, a=1a=-1, $b=1$, $\gamma_{1}=0.001$ and $\gamma_{2}=4$.
 We first train the baseline PINN model, which produces a large relative $L^{2}$ error of $0.511$ and relative $L^{1}$ error of $0.328$, demonstrating its inadequacy in resolving the sharp interface dynamics (Table 2).
@@
 Interestingly, unlike the previous case, there is minimal performance difference between the schemes using $IC1$ and $IC2$, suggesting that the impact of initial condition decay becomes less significant when strong oscillations are already present at the beginning.
-Figure 3: Exact solution of the Allen-Cahn Equation (Case II) with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 4: Best solutions of the Allen-Cahn Equation (Case II) $\gamma_{2}=4$.
-The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$.
-The hyper-parameter configuration is specified in Section 3.
-3.3 Cahn-Hilliard Equation
-Figure 5: Exact solution of the 1D Cahn-Hilliard Equation with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $9.75\times 10^{-4}$ and the relative $L^{1}$ error is $5.65\times 10^{-4}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 6: Best solutions of the 1D Cahn-Hilliard Equation.
-The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$.
-The hyper-parameter configuration is specified in Section 3.
-The Cahn-Hilliard equation, a variant of the Allen-Cahn equation, also models phase separation.
-This process describes how the components of a binary fluid spontaneously separate into distinct domains, each composed of a single component cahn1958free ; kim2016basic ; lee2014physical ; miranville2017cahn .
-The equation is expressed as:
-$u_{t}=\epsilon_{1}(-u_{xx}-\epsilon_{2}u_{xxxx}+(u^{3})_{xx}),\quad(t,x)\in(0,T]\times(-L,L),$ (5)
-$u(0,x)=u_{0}(x),\quad x\in[-L,L],$
-$u(t,-L)=u(t,L),\quad t\in[0,T],$
-where $\epsilon_{1}=10^{-2},\epsilon_{2}=10^{-4},T=1,L=1$.
-We specify the initial condition as $u_{0}(x)=-\cos(2\pi x)$.
-This PDE involves higher-order derivatives, making it more challenging to solve compared to the Allen-Cahn equation.
-For the ablation study in Section 2, the neural network is trained using the same hyperparameters outlined in Section 3.
-In this experiment, the full scheme yields the best results.
-Additionally, methods using initial condition IC2 outperform those with IC1.
-Except for the full schemes and the experiment with periodic neural networks for boundary conditions and initial condition IC2, none of the other approaches produce accurate solutions.
-The errors for all experiments are summarized in Table 2.
-The best results are illustrated in Figures 5 and 6.
-Table 2: Cahn-Hilliard Equation: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
-Ablation Settings Performance
-PBC IC1 (C=1) IC2 (C=0.1) SA Rel. $L^{2}$ error Rel. $L^{1}$ error
-$1.06\times 10^{-3}$ $5.61\times 10^{-4}$
-$9.75\times 10^{-4}$ $5.65\times 10^{-4}$
-$2.50\times 10^{-1}$ $5.87\times 10^{-2}$
-$3.87\times 10^{-3}$ $1.91\times 10^{-3}$
-$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
-$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
-$3.48\times 10^{-1}$ $2.94\times 10^{-1}$
-$2.37\times 10^{-1}$ $1.09\times 10^{-1}$
-$4.07\times 10^{-1}$ $3.45\times 10^{-1}$
-$1.55\times 10^{-2}$ $8.20\times 10^{-3}$
-$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
-$9.99\times 10^{-1}$ $9.98\times 10^{-1}$
-3.4 Kuramoto-Sivashinsky Equation
-Kuramoto–Sivashinsky equation is a fourth-order nonlinear partial differential equation, known for its chaotic behavior kudryashov1990exact ; michelson1986steady .
+Figure 4: Best solution: Allen-Cahn.
+B.3 Kuramoto-Sivashinsky PDE
+Kuramoto–Sivashinsky (KS) equation is a fourth-order nonlinear partial differential equation, known for its chaotic behavior Kudryashov (1990); Michelson (1986).
 It has served as a pivotal model in the study of complex dynamical phenomena observed in various physical systems.
-Consequently, the Kuramoto–Sivashinsky equation provides an excellent case study to demonstrate how PINNs can effectively simulate chaotic dynamics.
+Consequently, the KS equation provides an excellent case study to demonstrate how PINNs can effectively simulate chaotic dynamics.
 The equation is given by:
 $u_{t}=-u_{xx}-u_{xxxx}-uu_{x},\quad(t,x)\in(0,T]\times(a,b),$ (6)
-$u(0,x)=u_{0}(x),\quad x\in[a,b],$
-$u(t,a)=u(t,b),\quad t\in[0,T],$
-where $T=20,a=0,b=32\pi$, and $u_{0}(x)=\cos\frac{x}{16}(1+\sin\frac{x-1}{16})$. The parameters are the same as those mentioned in Section 3. The best results are shown in Figure 7 and 8:
-Figure 7: Exact solution of the 1D Kuramoto-Sivashinsky Equation with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $1.46\times 10^{-3}$ and the relative $L^{1}$ error is $1.02\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 8: Best solutions of the 1D Kuramoto-Sivashinsky Equation.
-The resulting relative $L^{2}$ error is $1.46\times 10^{-3}$ and the relative $L^{1}$ error is $1.02\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
-Table 3: Kuramoto-Sivashinsky Equation: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
+$u(0,x)=\cos(\frac{x}{16})(1+\sin\frac{x-1}{16}),\quad x\in[a,b],\quad u(t,a)=u(t,b),\,t\in[0,T],$
+where $T=20,a=0,b=32\pi$. The best results are shown in 5:
+Figure 5: Best solution: Kuramoto-Sivashinsky.
 Ablation Settings Performance
-PBC IC1 (C=1) IC2 (C=0.1) SA Rel. $L^{2}$ error Rel. $L^{1}$ error
+PBC IC1 IC2 SA Rel. $L^{2}$ error Rel. $L^{1}$ error
 $4.04\times 10^{-2}$ $1.46\times 10^{-2}$
-$1.46\times 10^{-3}$ $1.02\times 10^{-3}$
+$5.37\times 10^{-4}$ $1.02\times 10^{-3}$
 $1.68\times 10^{-2}$ $8.94\times 10^{-3}$
@@
 $3.01\times 10^{-1}$ $1.97\times 10^{-1}$
+Table 3: Kuramoto-Sivashinsky: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
 As in Table 3, solving this PDE system using a baseline PINN alone proves to be exceptionally challenging.
 The incorporation of periodic neural networks for boundary conditions, paired with initial condition IC2, consistently yields superior results compared to IC1.
-This suggests that, in contrast to the Allen-Cahn equation discussed in Section 3.1, the choice of initial condition plays a more significant role in the training process for this more complex system.
+This suggests that, in contrast to the Allen-Cahn equation discussed in Section B.2, the choice of initial condition plays a more significant role in the training process for this more complex system.
 This observation underscores the importance of carefully selecting both the boundary conditions and initial conditions when training neural networks for chaotic or higher-order PDE systems, where these factors contribute notably to the accuracy and stability of the solution.
-3.5 Gray-Scott Equation
-Reaction and diffusion of chemical species can produce a variety of patterns de1991turing ; nishiura1999skeleton , reminiscent of those often seen in nature.
-The Gray-Scott type system is one of classical mathematical models for chemical reactions gray1983autocatalytic ; gray1984autocatalytic ; liang2022reversible .
+B.4 Cahn-Hilliard Equation
+The Cahn-Hilliard equation, a variant of the Allen-Cahn equation, also models phase separation.
+This process describes how the components of a binary fluid spontaneously separate into distinct domains, each composed of a single component Cahn & Hilliard (1958); Kim et al.
+(2016); Lee et al.
+(2014); Miranville (2017).
+The equation is expressed as:
+$u_{t}=\epsilon_{1}(-u_{xx}-\epsilon_{2}u_{xxxx}+(u^{3})_{xx}),\quad(t,x)\in(0,T]\times(-L,L),$ (7)
+$u(0,x)=u_{0}(x),\quad x\in[-L,L],$
+$u(t,-L)=u(t,L),\quad t\in[0,T],$
+where $\epsilon_{1}=10^{-2},\epsilon_{2}=10^{-4},T=1,L=1$.
+We specify the initial condition as $u_{0}(x)=-\cos(2\pi x)$.
+This PDE involves higher-order derivatives, making it more challenging to solve compared to the Allen-Cahn equation.
+Figure 6: Best solution of the Cahn-Hilliard. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$.
+In this experiment, the full scheme yields the best results.
+Additionally, methods using initial condition IC2 outperform those with IC1.
+Except for the full schemes and the experiment with periodic neural networks for boundary conditions and initial condition IC2, none of the other approaches produce accurate solutions.
+The errors for all experiments are summarized in Table 4.
+The best results are illustrated in Figure 6.
+Ablation Settings Performance
+PBC IC1 IC2 SA Rel. $L^{2}$ error Rel. $L^{1}$ error
+$1.06\times 10^{-3}$ $5.61\times 10^{-4}$
+$9.75\times 10^{-4}$ $5.65\times 10^{-4}$
+$2.50\times 10^{-1}$ $5.87\times 10^{-2}$
+$3.87\times 10^{-3}$ $1.91\times 10^{-3}$
+$99\times 10^{-1} 9.9$$99\times 10^{-1}$
+$99\times 10^{-1} 9.9$$99\times 10^{-1}$
+$3.48\times 10^{-1}$ $2.94\times 10^{-1}$
+$2.37\times 10^{-1}$ $1.09\times 10^{-1}$
+$4.07\times 10^{-1}$ $3.45\times 10^{-1}$
+$1.55\times 10^{-2}$ $8.20\times 10^{-3}$
+$99\times 10^{-1} 9.9$$99\times 10^{-1}$
+$99\times 10^{-1} 9.9$$98\times 10^{-1}$
+Table 4: Cahn-Hilliard Equation: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
+B.5 Gray-Scott Equation
+Reaction and diffusion of chemical species can produce a variety of patterns (De Kepper et al., 1991; Nishiura & Ueyama, 1999), reminiscent of those often seen in nature.
+The Gray-Scott type system is one of classical mathematical models for chemical reactions (Gray & Scott, 1983, 1984; Liang et al., 2022).
 The general irreversible Gray-Scott equations describe such reactions:
@@
 This system is defined by two equations that describe the dynamics of two reacting substances:
-$u_{t}=\epsilon_{1}u_{xx}+b(1-u)-uv^{2},\quad(t,x)\in(0,T]\times(-L,L),$ (8)
-$v_{t}=\epsilon_{2}v_{xx}-(b+k)v+uv^{2},$
-$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad\forall x\in[-L,L],$
-$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad\forall t\in[0,T],$
-where $T=20,L=50$, $\epsilon_{1}=1,\epsilon_{2}=0.01$ are diffusion rates, $b=0.02$ is the "feeding rate" that adds $U$, $k=0.0562$ is the "killing rate" that removes $V$.
+$u_{t}=\epsilon_{1}u_{xx}+b(1-u)-uv^{2},$ (9)
+$v_{t}=\epsilon_{2}v_{xx}-(b+k)v+uv^{2},\quad(t,x)\in(0,T]\times(-L,L),$
+$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad x\in[-L,L],$
+$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad t\in[0,T],$
+where $T=20,L=50$, $\epsilon_{1}=1,\epsilon_{2}=0.01$ are diffusion rates, $b=0.02$ is the ”feeding rate” that adds $U$, $k=0.0562$ is the ”killing rate” that removes $V$.
 We set our initial conditions as:
-$u_{0}(x)=1-\frac{\sin(\pi(x-50)/100)^{4}}{2},\quad v_{0}(x)=\frac{\sin(\pi(x-50)/100)^{4}}{4}.$
+$u_{0}(x)=1-\frac{\sin(\pi(x-50)/100)^{4}}{2},$
+$v_{0}(x)=\frac{\sin(\pi(x-50)/100)^{4}}{4}.$
 In this ablation study, we not only evaluate the performance of self-adaptive neural networks but also test our scheme with mini-batching on the Gray-Scott equations, which model the densities of two interacting species.
@@
 These results suggest that IC2 is more effective for larger time-scale PDEs, as IC1 causes the initial condition to decay too quickly, reducing its influence during training.
-The predicted solutions from our best model are visualized in Figures 9 and 10, respectively.
-Figure 9: Best solutions of the 1D Gray-Scott Equation (u species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
-Figure 10: Best solutions of the 1D Gray-Scott Equation (v species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
-Figure 11 and Figure 12 show the comparison of the best prediction against the reference solution. Note that the majority of this error is attributed to the sharpness of the PDEs.
-Figure 11: Exact solution of the 1D Gray-Scott Eqaution (v specie) with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 12: Exact solution of the 1D Gray-Scott Eqaution (v specie) with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$.
-The hyper-parameter configuration is specified in Section 3.
-Table 4: Gray-Scott Equation: Relative $L^{2}$ errors of $u$ and $v$ equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
+The predicted solutions from our best model are visualized in Figures 7 and 8, respectively.
+Figure 7: Best solution of the Gray-Scott (u species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$.
+Figure 8: Best solution of the Gray-Scott (v species). The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$.
 Ablation Settings Performance
-PBC IC1 IC2 SA mini-batch Rel. $L^{2}$ error (u - eq.) Rel. $L^{2}$ error (v - eq.)
+PBC IC1 IC2 SA mini-batch $\mathcal{E}_{2}:u$ $\mathcal{E}_{2}:v$
 $8.98\times 10^{-4}$ $7.92\times 10^{-3}$
@@
 $9.21\times 10^{-2}$ $2.98\times 10^{-1}$
-3.6 Belousov-Zhabotinsky Equation
-A Belousov–Zhabotinsky reaction is one of a class of reactions that serve as a classical example of non-equilibrium thermodynamics.
-It typically involves the oxidation of organic compounds by bromine in an acidic medium zhabotinsky2007belousov ; cassani2021belousov .
-The system of PDEs takes the following form:
-$u_{t}=\epsilon_{1}u_{xx}+u+v-uv-u^{2},$ (9)
-$v_{t}=\epsilon_{2}v_{xx}+w-v-uv,$
-$w_{t}=\epsilon_{1}w_{xx}+u-w,\quad(t,x)\in(0,T]\times[-L,L]$
-$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad w(0,x)=w_{0}(x),\quad\forall x\in[-L,L]$
-$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad w(t,-L)=w(t,L),\quad\forall t\in[0,T],$
-where $T=2,L=1$, $\epsilon_{1}=10^{-5},\epsilon_{2}=2\times 10^{-5}$ are diffusion rates. We let the initial conditions be:
-$u_{0}(x)=\exp(-100(x+0.5)^{2}),\quad v_{0}(x)=\exp(-100x^{2}),\quad w_{0}(x)=\exp(-100(x-0.5)^{2}).$
-The system of PDEs involving three species provides a valuable opportunity to evaluate our proposed schemes using the techniques described in Section 2 for simulating multispecies dynamics.
-The results, summarized in Table 5, show that the full scheme consistently achieves the best overall performance, in line with previous experiments.
-Interestingly, the errors observed for methods using IC1 and those using IC2 are not significantly different.
-Both IC1 and IC2, when used independently, are capable of producing satisfactory results.
-However, when periodic neural networks are used to enforce boundary conditions, the performance declines and becomes less stable, even though self-adaptive training and mini-batching provide some improvements.
-This observation suggests that incorporating boundary conditions introduces additional complexity into the training process.
-Further investigation into their effects would be a promising direction for future research.
-Figure 13: Best solutions of the 1D Belousov-Zhabotinsky equation (u specie). The resulting relative $L^{2}$ error is $3.49\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
-Figure 14: Best solutions of the 1D Belousov-Zhabotinsky equation (v specie). The resulting relative $L^{2}$ error is $2.45\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
-Figure 15: Best solutions of the 1D Belousov-Zhabotinsky equation (w specie). The resulting relative $L^{2}$ error is $2.50\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
-The comparisons between the exact solutions of $(u,v,w)$ equations with corresponding best neural network predictions on the whole $(x,t)$ domain are shown in Figures 16, 17 and 18.
-Although there is a slight misalignment between each two of them, the model predictions closely match the corresponding numerical estimations.
-Figure 16: Exact solution of the 1D Belousov-Zhabotinsky equation (u specie) with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $3.49\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 17: Exact solution of the 1D Belousov-Zhabotinsky equation (v specie) Equation with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $2.45\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 18: Exact solution of the 1D Belousov-Zhabotinsky equation (w specie) Equation with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $2.50\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
-Table 5: Belousov-Zhabotinsky Equation: Relative $L^{2}$ errors of $u$, $v$ and $w$ equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
-Ablation Settings Performance: relative $L^{2}$ error
-PBC IC1 IC2 SA mini-batch u-equation v-equation w-equation
-$3.49\times 10^{-3}$ $2.45\times 10^{-3}$ $2.50\times 10^{-3}$
-$6.06\times 10^{-3}$ $4.47\times 10^{-3}$ $4.30\times 10^{-3}$
-$7.08\times 10^{-3}$ $4.77\times 10^{-3}$ $4.24\times 10^{-3}$
-$5.03\times 10^{-3}$ $3.10\times 10^{-3}$ $3.71\times 10^{-3}$
-$9.17\times 10^{-3}$ $9.29\times 10^{-3}$ $8.89\times 10^{-3}$
-$1.05\times 10^{-1}$ $1.78\times 10^{-2}$ $1.29\times 10^{-2}$
-$2.22\times 10^{-2}$ $1.33\times 10^{-2}$ $1.52\times 10^{-2}$
-$1.54\times 10^{-2}$ $9.17\times 10^{-3}$ $1.07\times 10^{-2}$
-$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
-$7.99\times 10^{-3}$ $5.05\times 10^{-3}$ $5.16\times 10^{-3}$
-$9.39\times 10^{-3}$ $9.41\times 10^{-3}$ $9.85\times 10^{-3}$
-$3.98\times 10^{-2}$ $2.15\times 10^{-2}$ $2.79\times 10^{-2}$
-$4.48\times 10^{-2}$ $2.71\times 10^{-2}$ $3.22\times 10^{-2}$
-$7.87\times 10^{-2}$ $3.95\times 10^{-2}$ $5.48\times 10^{-2}$
-3.7 Nonlinear Schröedinger Equation
-In theoretical physics, nonlinear Schröedinger equation is a nonlinear PDE, applicable to classical and quantum mechanics kato1987nonlinear ; kevrekidis2001discrete ; fibich2015nonlinear .
+Table 5: Gray-Scott Equation: Relative $L^{2}$ errors of $u$ and $v$ equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
+B.6 Nonlinear Schröedinger Equation
+In theoretical physics, nonlinear Schröedinger equation is a nonlinear PDE, applicable to classical and quantum mechanics (Kato, 1987; Kevrekidis et al., 2001; Fibich, 2015).
 The dimensionless equation of the classical field is
@@
 Using the same parameter settings and training steps with prior experiments, we obtain the following results:
-Figure 19: Exact solution of the real part of 1D Nonlinear Schroedinger Equation with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $1.80\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 20: Best real-part solutions of the 1D Nonlinear Schroedinger Equation.
-The resulting relative $L^{2}$ error is $1.80\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 21: Exact solution of the imaginary part of 1D Nonlinear Schroedinger Equation with the corresponding best network prediction and the absolute error difference.
-The resulting relative $L^{2}$ error is $2.58\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
-Figure 22: Best imaginary-part solutions of the 1D Nonlinear Schroedinger Equation.
-The resulting relative $L^{2}$ error is $2.58\times 10^{-3}$.
-The hyper-parameter configuration is specified in Section 3.
+Figure 9: Best real-part solutions of the Nonlinear Schröedinger Equation.
+Figure 10: Best imaginary-part solutions of the Nonlinear Schröedinger Equation.
 Table 6 presents the results of all experiments.
@@
 Moreover, without the use of self-adaptive neural networks or mini-batching, accurately obtaining the solution becomes challenging, highlighting the importance of both methods in ensuring more stable training.
-Table 6: Nonlinear Schroedinger Equation: Relative $L^{2}$ errors of real and imaginary equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
-Ablation Settings Performance: relative $L^{2}$ error
-PBC IC1 IC2 SA mini-batch Real part Imaginary part
-$1.80\times 10^{-3}$ $2.58\times 10^{-3}$
-$1.84\times 10^{-3}$ $3.27\times 10^{-3}$
-$7.34\times 10^{-2}$ $1.23\times 10^{-1}$
-$8.32\times 10^{-2}$ $1.39\times 10^{-1}$
-$1.10\times 10^{-2}$ $2.10\times 10^{-2}$
-$1.70\times 10^{-2}$ $3.27\times 10^{-2}$
+Ablation Settings Performance: $\mathcal{E}_{2}$
+PBC IC1 IC2 SA mini-batch Real part Imag part
+$1.80\times 10^{-4}$ $2.58\times 10^{-4}$
+$1.84\times 10^{-4}$ $3.27\times 10^{-4}$
+$7.34\times 10^{-3}$ $1.23\times 10^{-2}$
+$8.32\times 10^{-3}$ $1.39\times 10^{-2}$
+$1.10\times 10^{-3}$ $2.10\times 10^{-3}$
+$1.70\times 10^{-3}$ $3.27\times 10^{-3}$
 $1.00\times 10^{-1}$ $1.83\times 10^{-1}$
@@
 $3.63\times 10^{-1}$ $5.06\times 10^{-1}$
-4 Conclusion
-We provide an ablation study of hard-constraints and self-adaptive loss weights to study the solution dependency of time-dependent PDEs on its initial conditions.
-The transformation provided by hard-constraints significantly alleviates the training challenges typically associated with stiff time-dependent PDEs subject.
-Our study demonstrates effectiveness across seven distinct types of such PDEs, encompassing scalar-valued outputs, vector-valued outputs, and even complex-valued solutions.
-This approach not only enhances training efficiency but also improves the accuracy and stability of the learned solutions, particularly in the presence of stiffness, where conventional PINNs often struggle to converge.
-Although we provide an ablation study of hard-constraints and self-adaptive loss weights, our list of possible training remedies for PINNs is not exhaustive.
-Moreover, our tests should also include elliptical PDEs with high oscillatory coefficients.
-Future work includes extending the hard-cosntraint framework to two- and three-dimensional PDEs, which pose additional challenges in both architecture design and computational cost.
-Ongoing research also focuses on identifying optimal transformations $\psi$ and $\phi$, which govern the embedding of IC and BC into the network structure, with the goal of further enhancing performance and generalizability.
-Acknowledgements
-BH and MZ thank Prof.
-Yannis Kevrekids (JHU) for the discussion of well-conditioning of PDE solution maps, and Prof.
-George Karniadakis (Brown) for his suggestion on training enhancement methods for PINNs.
-BH developed and implemented the algorithm, ran the simulation and analyzed the data.
-UB and MZ designed the reserach.
-All authors wrote the paper.
-MZ is supported by NSF-AoF grant #$\#2225507$ and Ralph E.
-Powe Junior Faculty Enhancement Award from ORAU for FY24242424.
-Supplementary Material
-The source code, initial training data, initial weights, Chebfun code, and ground truth solutions used for benchmarking will provided in https://github.com/baolihao/Training-PINNs-with-Hard-Constraints-and-Adaptive-Weights.
-The list of hyperparameters and computational resources used to produce the results will be discussed later in this appendix.
+Table 6: Nonlinear Schröedinger Equation: Relative $L^{2}$ errors of real and imaginary equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
+B.7 Linear Advection
+The linear advection equation models the transport of a quantity by a constant velocity field and serves as a fundamental benchmark for testing numerical methods due to its simplicity and rich structure LeVeque (2002); Toro (2013).
+It describes the propagation of a scalar profile without deformation and is given by:
+$u_{t}+cu_{x}=0,\quad\quad\quad(t,x)\in(0,T]\times(0,2\pi),$ (11)
+$u(0,x)=\sin(x),\quad x\in[0,2\pi],$
+$u(t,0)=u(t,2\pi),\quad t\in[0,T].$
+Unlike dissipative PDEs such as the Allen–Cahn or Cahn–Hilliard equations, the linear advection equation conserves the shape of the initial profile as it translates.
+This problem becomes “stiffer” as the advection velocity increases.
+Theorem 3.
+Let $u(t,x)$ denote the solution to the linear advection equation (11).
+With the transformed neural approximation in (3) and $\psi$ and $\eta$ satisfying the conditions in (4), then $u(t,x)$ satisfies the PDE iff $\eta$ satisfies the forced linear transport equation:
+$\eta_{t}+c\eta_{x}=-(\psi_{t}+c\psi_{x}).$
+This PDE is well-posed and stable for PINN training provided the following conditions hold:
+  • The function $f(t,x):=-(\psi_{t}+c\psi_{x})$ satisfies $f\in C^{1}([0,T]\times\Omega)$.
+  • If $f(t,x)$ is uniformly bounded or decays in time, then the solution $\eta$ remains stable and avoids growth-induced overfitting.
+For our choice, we have $\psi(t,x)=u_{0}(x)e^{-\alpha t}$ and $\phi(t,x)=t$ with $\alpha>0$, ensuring $\eta(0,x)=0$.
+Then, the forcing term satisfies $\eta_{t}+c\eta_{x}=(\alpha\sin{x}-c\cos{x})e^{-\alpha t}$.
+Since this forcing is smooth and decays exponentially in time, it guarantees both the stability and regularity of the neural network $\eta(t,x)$.
+Here, we compare the results produced by our method with those from a baseline PINN under the same hyperparameter settings.
+In our experiments, both models were able to converge at low velocities; however, we observed that higher advection velocities introduced stiffness.
+Our method exhibited greater robustness and was less affected by stiffness compared to the baseline PINN.
+Figure 11: Periodic advection results after 50,000 iterations.
+The implementation details are the same as in Section B.1, except that we perform $10$ independent replications of the experiment to account for the randomness in neural network weight initialization, rather than using the same initial weights.
+Table 7 shows the averaged relative $L_{1}$$L_{2}$-error over $10$ independent repetitions.
+As the advection velocity increases and the PDE becomes stiffer, making it harder for baseline PINN to train.
+However, our method consistently produces better results.
+This trend is clearly reflected in Table 7.
+Even at low velocities, our method demonstrates significantly better accuracy compared to the baseline PINN.
+As $c$ increases, both the relative $L_{1}$ and $L_{2}$ errors grow, however, the errors associated with the PINN rise significantly faster than those of our method.
+The most striking example occurs at $c=50$, where the base line PINN fails to approximate the solution, while our method still produces stable and good results.
+This behavior is evident in Table 7 and Figure 11.
+Table 8 shows the ablation study of linear advection equation.
+Solver $c$ Rel. $L_{1}$ error Rel. $L_{2}$ error
+30 $0.0030$ $0.0041$
+HC-PINN 40 $0.0141$ $0.0159$
+50 $0.0147$ $0.0164$
+60 $0.0171$ $0.0182$
+30 $0.0478$ $0.0579$
+CINN(Braga-Neto, 2022) 40 $0.0714$ $0.0852$
+50 $0.4692$ $0.5365$
+60 $0.7293$ $0.7134$
+30 $0.0873$ $0.1003$
+Baseline 40 $0.3918$ $0.4395$
+PINN 50 $0.7140$ $0.7797$
+60 $0.8339$ $0.8664$
+Table 7: Comparison of HC-PINN, CINN and baseline PINN across different advection speeds $c$ over 10 independent repetitions.
+Ablation Settings Performance
+PBC IC SA Rel. $L^{2}$ error Rel. $L^{1}$ error
+$1.82\times 10^{-2}$ $1.03\times 10^{-2}$
+$9.75\times 10^{-2}$ $8.65\times 10^{-2}$
+$7.50\times 10^{-1}$ $7.87\times 10^{-1}$
+$3.87\times 10^{-2}$ $5.47\times 10^{-2}$
+$7.44\times 10^{-1}$ $8.01\times 10^{-1}$
+$07\times 10^{-1} 1.1$$19\times 10^{-1}$
+$65\times 10^{-1} 7.7$$71\times 10^{-1}$
+$8.66\times 10^{-1}$ $8.34\times 10^{-1}$
+Table 8: Linear Advection Equation (c=60): Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.